### v1

# Orbit structures and complexity in Schubert and Richardson Varieties

Abstract.
The goal of this paper is twofold. Firstly, we provide a type-uniform formula for the torus complexity of the usual torus action on a Richardson variety, by developing the notion of algebraic dimensions of Bruhat intervals, strengthening a type A result by Donten-Bury, Escobar and Portakal. In the process, we give an explicit description of the torus action on any Deodhar component as well as describe the root subgroups that comprise the component. Secondly, we exhibit a codimension preserving bijection between Levi-Borel subgroup (a Borel subgroup of a Levi subgroup in a reductive algebraic group) orbits in the big open cell of a Schubert variety and torus orbits in the big open cell of a distinguished Schubert subvariety. This bijection has many applications: a type-uniform formula for the Levi-Borel complexity of the usual Levi-Borel subgroup action on a Schubert variety as well as an indexing set for Levi-Borel subgroup orbits in a Schubert variety. We conclude by extending the Levi-Borel complexity results to a large class of Schubert varieties in the partial flag variety.
1. Introduction
1.1. Group orbits in the full flag variety
Let $G$ be a complex, connected, reductive algebraic group of rank $r$. Fix a maximal torus $T$ in $G$, and a Borel subgroup $B$ of $G$ containing $T$. The homogeneous space $G/B$ is a smooth projective variety known as the full flag variety. The study of the flag variety first arose out of need to formalize and justify the enumerative geometry of H. Schubert as laid out in Hilbert’s 15th problem. These varieties are a central theme in much of the mathematics of the 20th century and beyond, with deep and fundamental connections to algebraic geometry, Lie theory, representation theory, algebraic combinatorics, and commutative algebra.
Of particular importance has been the study of orbits and orbit closures in the flag variety. Famously, the $B$-orbits for the action of $B$ by left translation yield a cellular filtration of $G/B$; these orbits, denoted $X^{\circ}_{w}$, are referred to as Schubert cells and are indexed by elements $w$ of the Weyl group $W$ of $G$. In his seminal work [C94], C. Chevalley introduced the now ubiquitous Bruhat order to describe the inclusion order of $B$-orbit closures in $G/B$. These orbit closures, or as they are more commonly known, Schubert varieties $X_{w}$ for $w\in W$, are well-studied varieties that boast a rich combinatorial structure that encodes many facets of their geometry. They are known to be normal, Cohen–Macaulay, and have rational singularities [DL81, RR85, R85].
A stratification of $G/B$ via the orbits of the opposite Borel subgroup $B^{-}$ also exists; the closure of these orbits are the opposite Schubert varieties $X^{w}$ for $w\in W$. The Richardson variety $\mathcal{R}_{u,v}$ for $u,v\in W$ is defined to be the intersection $\mathcal{R}_{u,v}:=X_{v}\cap X^{u}$. Though Richardson varieties have arisen in the literature in a number of guises, initially in the work of W. V. D. Hodge [H42], they were first introduced in their modern form by D. Kazhdan and G. Lusztig [KL80]. They were first shown to be irreducible by Deodhar [D85]. Later, R. W. Richardson, in whose honor the varieties are named, studied more general intersections of double cosets and their closures, proving that they are reduced and irreducible [R92]. M. Brion proved that Richardson varieties are normal, Cohen–Macaulay, and have rational singularities [B02].
While the Borel subgroup orbits are perhaps the most fundamental class of orbits in the full flag variety, they are not the only orbits that have been explored. The orbits of $T$ in the flag variety were first studied by Klyachko [K85]. In related work, I. M. Gelfand and V. V. Serganova proved that realizable matroids correspond to torus orbits in the Grassmannian (a partial flag variety) [GS87]. This connection between matroid theory can be extended to torus orbits in the full flag variety via flag matroids [CDMS22, §6.5].
1.2. Torus orbits in Schubert and Richardson Varieties
It is also instructive to study group orbits and their closures within Schubert and Richardson varieties. Of particular interest in this setting is the statistic on the set of orbits of a reductive algebraic group $H$ known as the $H$-complexity, a notion which we now make precise.
If an algebraic group $H$ acts on a variety $X$ by a morphism of algebraic varieties, we say that $X$ is an $H$-variety. We denote the set of $H$-orbits in $X$ by $\mathcal{O}_{H}(X)$.
Definition 1.1.
Let $H$ be a reductive algebraic group and $B_{H}$ a Borel subgroup of $H$. The $H$-complexity of an $H$-variety $X$, denoted $c_{H}(X)$, is the minimum codimension of a $B_{H}$-orbit in $X$.
The normal $H$-varieties with $H$-complexity equal to 00 are the $H$-spherical varieties. Spherical varieties generalize several important classes of algebraic varieties including toric varieties, projective rational homogeneous spaces and symmetric varieties. The birational models of a given spherical variety are classified via Luna-Vust theory using colored fans [LV83, K91], a generalization of the fans which classify toric varieties. The birational classes themselves are classified by the spherical systems of Luna [L01].
Unless otherwise stated, in this paper the action of any subgroup of $G$ on any subvariety of $G/B$ will always be left translation. We will refer to this as the usual action.
For any $w\in W$, both $X_{w}$ and $X^{w}$ are $T$-varieties for the usual action. We denote the Bruhat order on $W$ by $\leq$, writing $u\leq v$ if and only if $X_{u}\subseteq X_{v}$. For $u\leq v\in W$, $\mathcal{R}_{u,v}$ is nonempty and is a normal $T$-variety for the usual action. The maximal torus $T$ is a reductive algebraic group whose only Borel subgroup is $T$ itself. Hence, a normal $T$-variety with $T$-complexity equal to 00 is a $T$-spherical variety (that is, a toric variety).
P. Karuppuchamy provided a succinct, root-system uniform classification of the toric Schubert varieties [K13]. E. Tsukerman and L. Williams introduced Bruhat interval polytopes to study the geometry of Richardson varieties and noted connections to torus orbits [TW15]. This connection was expanded on by E. Lee, M. Masuda, and S. Park, yielding a classification of toric Richardson varieties in $G/B$ where $G$ is of Dynkin-type $A$ [LMP21a], with further results by C. Gaetz [G23] . In [LMP21b], the same authors give a classification of $T$-complexity 1 Schubert varieties in $G/B$ when $G$ is of Dynkin-type $A$. A classification of toric Richardson varieties in any full flag variety was provided by M. Can and P. Saha in [CS23b].
Our first main result continues this line of research, giving a root-system uniform combinatorial formula for the $T$-complexity of a Richardson variety in any full flag variety. We do so by defining and developing the notion of algebraic dimensions for arbitrary Bruhat intervals $[u,v]$ in Section 3, where we provide multiple formulas for $\operatorname{ad}(u,v)$.
Theorem 1.2.
For $u\leq v\in W$, the $T$-complexity of the Richardson variety equals
$c_{T}(\mathcal{R}_{u,v})=\ell(v)-\ell(u)-\operatorname{ad}(u,v).$
Moreover, we have that
$\operatorname{ad}(u,v)=\max_{u\leq w\leq v}\ell(v)-\ell(w)\text{ where }\mathcal{R}_{w,v}\text{ is toric}.$
One highlight of Theorem 1.2 is that our method works uniformly across all Lie types, and our key lemmas work uniformly across geometric realizations for any Coxeter groups, while the theory of Bruhat interval polytopes is well-developed in type $A$, but not as developed in other types. Moreover, Theorem 1.2 can be seen as a generalization of the main result of [DEP23] to other Lie types (with different arguments). Along the way, in Theorem 3.18, we are also able to explicitly describe the torus action on any Deodhar component by enumerating the root subgroups that comprise the component, strengthening an argument of [CS23b].
In the case of $u=\mathrm{id}$, Theorem 1.2 simplifies to a simple, type-uniform combinatorial formula for the torus complexity of a Schubert variety.
Corollary 1.3.
For $w\in W$, the $T$-complexity of the Schubert variety $X_{w}$ equals
$c_{T}(X_{w})=\ell(w)-\operatorname{supp}(w).$
1.3. Levi-Borel subgroup orbits in Schubert Varieties
We now turn our attention to the orbits of Levi-Borel subgroups (defined below) in a Schubert variety. Our work uncovers a deep structural relationship between the Levi-Borel orbits in the big cell of a Schubert variety and the $T$-orbits in the big cell of a distinguished Schubert subvariety. Specifically, we exhibit a codimension-preserving bijection between these sets of orbits.
Our choice of $T$ and $B$ determine the root system $\Phi$ and a set of simple roots $\Delta=\{\alpha_{1},\ldots,\alpha_{r}\}$, respectively. The Weyl group $W$ of $G$, is generated by the set of simple reflections $\{s_{i}\!:=\!s_{\alpha_{i}}|\alpha_{i}\!\in\!\Delta\}$. For $I\!\subseteq\!\Delta$, $W_{I}$ is the subgroup of $W$ generated by $\{s_{\alpha_{i}}\>|\>\!i\!\in\!I\}$.
The standard parabolic subgroups of $G$ containing $B$ are indexed by subsets of $\Delta$. The standard parabolic subgroup associated to $I$ is $P_{I}=BW_{I}B$ with Levi decomposition
$P_{I}=L_{I}\ltimes U_{I},$
where $L_{I}$ is a reductive subgroup called a Levi subgroup, and $U_{I}$ is the unipotent radical of $P_{I}$. Define $B_{L_{I}}:=L_{I}\cap B$. Then $B_{L_{I}}$ is a Borel subgroup of $L_{I}$, and we shall refer to such subgroups as Levi-Borel subgroups. Our second main result concerns the orbits of Levi-Borel subgroups in a Schubert cell.
Theorem 1.4.
Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$. The map $\mathfrak{O}\!:\!\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{w}})\!\rightarrow\!\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\!\mapsto\!B_{L_{I}}\prescript{I}{}{w}x$, where $x$ is any point in $\Theta$, is a codimension-preserving bijection. That is,
$\dim(X_{w}^{\circ})-\dim(\mathfrak{O}(\Theta))=\dim(X^{\circ}_{\prescript{I}{}{w}})-\dim(\Theta),$
for every $\Theta\in\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{w}})$.
Theorem 1.4 allows us to intertwine the study of Levi-Borel orbits with the vast literature on torus orbits and varieties, including our own Corollary 1.3, and has myriad applications which we now detail. Our first application is Proposition 4.9 which provides a formula for the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$. In general, the formula given in Proposition 4.9 is non-trivial to compute. However, in the case where the Levi subgroup $L_{I}$ acts on the Schubert variety $X_{w}$ the formula simplifies to a closed form.
While $B_{L_{I}}$ acts on any $X_{w}$, since $B_{L_{I}}\subseteq B$, the same is not true for $L_{I}$. The stabilizer of $X_{w}$ in $G$ for the usual action is the standard parabolic subgroup $P_{{\mathcal{D}}_{L}(w)}$ [BL00, Lemma 8.2.3], where ${\mathcal{D}}_{L}(w)$ is the left descent set of $w$ defined in Section 2. Thus, the Levi subgroups $L_{I}\leq P_{I}\leq P_{{\mathcal{D}}_{L}(w)}$ for $I\subseteq{\mathcal{D}}_{L}(w)$ are a family of reductive algebraic groups which act on $X_{w}$. Our third main result is a root-system uniform, combinatorial formula for the $L_{I}$-complexity of any Schubert variety which is an $L_{I}$-variety in any full flag variety.
Theorem 1.5.
Let $w\in W$ and suppose $L_{I}$ acts on the Schubert variety $X_{w}$ (equivalently, $I\subseteq\mathcal{D}_{L}(w)$). If $w={{}_{I}w}\prescript{I}{}{w}$ is the left parabolic decomposition of $w$ with respect to $I$, then
$c_{L_{I}}(X_{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
Theorem 1.5 is a refinement and generalization of numerous earlier results. The second author and A. Yong initiated a study of $L_{I}$-complexity 00 Schubert varieties in [HY22], and therein conjectured a classification in terms of spherical elements of a Coxeter group. Subsequently, both authors of this paper and A. Yong proved this conjecture first for $G$ of Dynkin-type $A$ [GHY23], and later for any full flag variety [GHY24] (the latter was proved contemporaneously via different methods by M. Can and P. Saha [CS23a]). Using work of R. S. Avdeev and A. V. Petukhov [AP14], the classification of $L_{I}$-complexity 00 Schubert varieties may be interpreted as a generalization of results of P. Magyar, J. Weyman, and A. Zelevinsky [MWZ99] and J. Stembridge [S01, S03] on spherical actions on (products of) flag varieties. The interested reader may see [HY22, Theorem 2.4] for the details.
As a second application of Theorem 1.4, we are able to describe the orbits of Levi-Borel subgroups in the Schubert variety itself.
Corollary 1.6.
Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$. There is a bijection between $\mathcal{O}_{B_{L_{I}}}(X_{w})$ and
$\bigsqcup_{u\leq w}\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})\times\{u\}$
We highlight that in the $I=\Delta$ case, $B_{L_{I}}=B$ and $\prescript{I}{}{u}=\mathrm{id}$ so each $\mathcal{O}_{T}(\prescript{I}{}{u})$ is a singleton set. Hence, in this case Corollary 1.6 recovers the Bruhat decomposition, namely that the Borel subgroup orbits in a Schubert variety are in bijection with the set of all $u\leq w$. In this way, Theorem 1.4 and Corollary 1.6 can be viewed as the first steps towards a framework that aims to generalize the study of Borel subgroup orbits in the full flag variety to the study of Levi-Borel subgroup orbits in Schubert varieties. Our next step, with work in progress, is understanding the inclusion order on these orbit closures. The combinatorial indexing set developed in this paper is an important cornerstone of this work. Throughout the years, the study of such Bruhat orders has led to much fruitful research. One related example, is that of the orbits of spherical Levi subgroups in Dynkin type $A$. These orbits are indexed by $(p,q)$-clans [MO90, Y97], and their inclusion order is described via clan combinatorics [W16].
As a final application of Theorem 1.4, in Theorem 4.13 we compute the $L_{I}$-complexity of a large class of Schubert varieties in the partial flag variety. This leads to Corollary 4.14, which is a root-system uniform, combinatorial classification of toric Schubert varieties in any partial flag variety; as far as the authors are aware this classification is new.
The outline of this paper is as follows. In Section 2, we introduce the necessary background, including Schubert geometry and Coxeter groups. In Section 3, we fully develop the notion of algebraic dimensions and establish Theorem 1.2. In Section 4, we prove Theorem 1.4, Theorem 1.5, and Corollary 1.6 with one of our main tools being an equivariant isomorphism that allows us to use results in Section 3. Section 4 concludes with an extension of the Borel-Levi subgroup complexity results to Schubert varieties in partial flag varieties.
2. Preliminaries
We first recall some preliminaries and set notation which will be useful throughout.
2.1. Combinatorics of the Weyl group
Consider the following data associated to elements of the Weyl group $W$. For $w\in W$, its Coxeter length is the smallest $\ell$ such that $w$ can be written as product of $\ell$ simple reflections. Such an expression is called a reduced word or a reduced expression of $w$.
[…]
The support of $w$ is
$\operatorname{Supp}(w)=\{\alpha_{i}\in\Delta\>|\>s_{i}\text{ appears in any/all reduced words of }w\}.$
The cardinality of $\operatorname{Supp}(w)$ is written as $\operatorname{supp}(w)=|\operatorname{Supp}(w)|$. For $w\in W$, its left descent set and the right descent set are
${\mathcal{D}}_{L}(w)=\{\alpha_{i}\in\Delta\>|\>\ell(s_{i}w)<\ell(w)\},\quad{\mathcal{D}}_{R}(w)=\{\alpha_{i}\in\Delta\>|\>\ell(ws_{i})<\ell(w)\},$
respectively. And its left inversion set and the right inversion set are
${\mathcal{I}}_{L}(w)=\{\alpha\in\Phi^{+}\>|\>w(\alpha)\in\Phi^{-}\},\quad{\mathcal{I}}_{R}(w)=\{\alpha\in\Phi^{+}\>|\>w^{-1}(\alpha)\in\Phi^{-}\},$
respectively, with $|\mathcal{I}_{L}(w)|=|\mathcal{I}_{R}(w)|=\ell(w)$. There are equivalent ways to write down the inversion sets. For example,
$\mathcal{I}_{L}(w)=\Phi^{+}\cap w(\Phi^{-})=\{\alpha\in\Phi^{+}\>|\>\ell(s_{\alpha}w)<\ell(w)\}.$
Given a length additive product $uv$, [B02, Ch. VI, §1, Cor. 2 of Prop. 17] proves that
(1) ${\mathcal{I}}_{L}(uv)={\mathcal{I}}_{L}(u)\sqcup u({\mathcal{I}}_{L}(v)).$
The (strong) Bruhat order is a partial order on $W$ that is generated by
$w<ws_{\alpha}\text{ for }\alpha\in\Phi^{+},\text{ if }\ell(w)<\ell(ws_{\alpha}).$
Write $[u,v]:=\{w\in W\>|\>u\leq w\leq v\}$ for a Bruhat interval. It is well-known (see [BB05, Lemma 2.7.3]) that any interval $[u,v]$ of rank $2$, i.e. $\ell(v)-\ell(u)=2$, has $4$ elements forming a “diamond”. Here are some properties of the Bruhat order. See [BB05, Section 2.2].
Proposition 2.1 (Subword property).
Fix a reduced word $w=s_{i_{1}}\cdots s_{i_{\ell}}$. Then $u\leq w$ in the Bruhat order if and only if there exists a subexpression of $s_{i_{1}}\cdots s_{i_{\ell}}$ that equals $u$.
Proposition 2.2 (Lifting property).
Suppose $u<w$ and $s\in\mathcal{D}_{R}(w)\setminus\mathcal{D}_{R}(u)$, then $u\leq ws$ and $us\leq w$.
2.2. Root subgroups
For $\alpha\in\Phi$, let $U_{\alpha}$ be the root subgroup corresponding to $\alpha$. Given $w\in W$, define $U_{w}:=U\cap wU^{-}w^{-1}$, where $U$ and $U^{-}$ are the unipotent parts of $B$ and $B^{-}:=w_{0}Bw_{0}$, respectively. A sequence of subgroups $H_{1},\ldots,H_{k}$ in an algebraic group $H$ is said to directly span $H$ if the product morphism $H_{1}\times\cdots\times H_{k}\rightarrow H$ is a bijection. The subgroup $U_{w}$ of $U$ is closed and normalized by $T$. Hence [B91, §14.4] implies that $U_{w}$ is directly spanned (in any order) by the root subgroups contained in it. Thus
$U_{w}:=U\cap wU^{-}w^{-1}=\left(\prod_{\alpha\in\Phi^{+}}U_{\alpha}\right)\bigcap\left(w\prod_{\alpha\in\Phi^{-}}U_{\alpha}w^{-1}\right)=\left(\prod_{\alpha\in\Phi^{+}}U_{\alpha}\right)\bigcap\left(\prod_{\alpha\in\Phi^{-}}U_{w(\alpha)}\right),$
where the second equality is [B91, §14.4] applied to $U$ and $U^{-}$, and the last equality is [J03, Part II, 1.4(5)]. Thus $U_{\alpha}$ is contained in $U_{w}$ if and only if $\alpha\in{\mathcal{I}}(w)$. Hence
(2) $U_{w}=\prod_{\alpha\in{\mathcal{I}}(w)}U_{\alpha}.$
Let $w=ad$ be the left parabolic decomposition of $w$ with respect to $I$, where $a\in W_{I}$ and $d\in\prescript{I}{}{W}$. Define $V_{d}:=aU_{d}a^{-1}$.
Lemma 2.3.
The group $V_{d}$ is a closed subgroup of $U_{w}$ normalized by $T$. Indeed, $U_{w}$ is directly spanned in any order by the subgroups $U_{a}$ and $V_{d}$, that is, $U_{w}=U_{a}V_{d}=V_{d}U_{a}$.
Proof.
The group $V_{d}$ is closed since $U_{d}$ is closed. Now, let $t\in T$. Since $a\in W$, we have $t^{-1}a=at^{\prime}$ for some $t^{\prime}\in T$. Hence $a^{-1}t=(t^{\prime})^{-1}a^{-1}$. These results, along with the fact that $U_{d}$ is normalized by $T$, yield
$t^{-1}V_{d}t=t^{-1}aU_{d}a^{-1}t=t^{-1}aU_{d}a^{-1}t=at^{\prime}U_{d}(t^{\prime})^{-1}a^{-1}=aU_{d}a^{-1}=V_{d},$
allowing us to conclude that $V_{d}$ is normalized by $T$. For the remaining claims,
$U_{w}=\prod_{\alpha\in{\mathcal{I}}(ad)}U_{\alpha}$ $=\left(\prod_{\alpha\in{\mathcal{I}}(a)}U_{\alpha}\right)\left(\prod_{\alpha\in a({\mathcal{I}}(d))}U_{\alpha}\right)$
$=U_{a}\left(\prod_{\alpha\in{\mathcal{I}}(d)}U_{a(\alpha)}\right)=U_{a}\left(aU_{d}a^{-1}\right)=U_{a}V_{d}$
where the first equality is (2), the second is (1) combined with the length additivity of the left parabolic decomposition, and the fourth is [J03, Part II, 1.4(5)] and (2). In the same way, one can show $U_{w}=V_{d}U_{a}$. We conclude that $U_{w}$ is directly spanned in any order by its subgroups $U_{a}$ and $V_{d}$. ∎
3. The torus complexity of Richardson varieties
In this section, we establish a formula for the complexity of the usual torus action on any Richardson variety $\mathcal{R}_{u,v}$ where $u\leq v$ in the Bruhat order.
Definition 3.1.
The (undirected) Bruhat graph on a Weyl group $W$ is the graph $\Gamma$ with vertex set $W$ and edges $w\sim s_{\alpha}w$ for $\alpha\in\Phi^{+}$. For each edge $w\sim s_{\alpha}w$, we say that it has label or weight $\alpha$, and write $\operatorname{wt}(w,s_{\alpha}w)=\alpha$. For $u\leq v$, let $\Gamma(u,v)$ be the Bruhat graph restricted to the vertex set $[u,v]$.
Definition 3.2.
For $u\leq v$, let $\operatorname{AD}(u,v)$ be the $\mathbb{R}$-span of all edge labels in $\Gamma(u,v)$, i.e.
$\operatorname{AD}(u,v)=\mathrm{span}_{\mathbb{R}}\{\operatorname{wt}(x,y)\>|\>u\leq x<y\leq v\}.$
Let $\operatorname{ad}(u,v)=\dim\operatorname{AD}(u,v)$ be the algebraic dimension of the Bruhat interval $[u,v]$.
In the following subsection 3.1, we fully develop the notion of algebraic dimension, and in subsection 3.2, we finish the proof of Theorem 1.2 by Deodhar decompositions.
3.1. The algebraic dimension of Bruhat intervals
We first discuss some useful tools.
Lemma 3.3.
[D91, Lemma 3.1] For $\alpha,\beta,\gamma,\tau\in\Phi^{+}$, if $s_{\alpha}s_{\beta}=s_{\gamma}s_{\tau}\neq 1$, then $\mathrm{span}_{\mathbb{R}}(\alpha,\beta)=\mathrm{span}_{\mathbb{R}}(\gamma,\tau)$ is $2$-dimensional.
We will be using this lemma multiple times, primarily on Bruhat intervals of rank $2$.
Lemma 3.4.
For $\alpha_{i}\in\Delta$ and $\alpha\in\Phi^{+}$, if $w$ covers both $ws_{i}\neq s_{\alpha}w$, then both $ws_{i}$ and $s_{\alpha}w$ cover $s_{\alpha}ws_{i}$. Dually, if $w$ is covered by both $ws_{i}\neq s_{\alpha}w$, then $s_{\alpha}ws_{i}$ covers $ws_{i}$ and $s_{\alpha}w$.
Proof.
For the first statement, it suffices to show that $s_{\alpha}w$ has a right descent at $\alpha_{i}$. If not, by the lifting property (Proposition 2.2) on $w>s_{\alpha}w$, we must have $ws_{i}\geq s_{\alpha}w$, which is a contradiction since they are different elements of the same length.
For the dual statement, it suffices to show that $s_{\alpha}w$ does not have a right descent at $\alpha_{i}$. Assume the opposite. By the first statement on $s_{\alpha}w$, $ws_{i}$ must be covered by both $w$ and $s_{\alpha}ws_{i}$, contradicting $w<ws_{i}$. ∎
We now describe multiple ways to simplify the computation of $\operatorname{AD}(u,v)$.
Lemma 3.5.
$\operatorname{AD}(u,v)$ is spanned by the labels of all cover relations incident to $u$ inside $[u,v]$, and dually, by all cover relations incident to $v$ inside $[u,v].$
Proof.
We first show that $\operatorname{AD}(u,v)$ is spanned by all labels of cover relations in $[u,v]$. We say that an edge $x\sim y$ in the Bruhat graph has length $|\ell(x)-\ell(y)|$ and it suffices to show that a label of length greater than $1$ can be linearly generated by edge labels with strictly smaller length. By Proposition 3.3 of [D91], for any edge $x_{1}\sim x_{6}$ such that $\ell(x_{1})<\ell(x_{6})-1$ in the Bruhat graph, there are $x_{2},x_{3},x_{4},x_{5}\in[x_{1},x_{6}]$ such that $x_{1}$ is connected to $x_{i}$, $x_{i}$ is connected to $x_{j}$ and $x_{j}$ is connected to $x_{6}$ for all $i\in\{2,3\}$ and $j\in\{4,5\}$. Applying Lemma 3.3 to the reflections corresponding to the edge labels of $x_{1}\sim x_{2}\sim x_{4}\sim x_{6}$ and $x_{1}\sim x_{6}$, we have $\operatorname{wt}(x_{1},x_{6})\in\mathrm{span}_{\mathbb{R}}\{\operatorname{wt}(x_{1},x_{2}),\operatorname{wt}(x_{2},x_{4})\}$.
Next, let $V$ be the vector space spanned by the labels of all cover relations incident to $u$, and we need to show that $\operatorname{wt}(x,y)\in V$ for all $x\lessdot y$ inside $[u,v]$. We use induction on $\ell(x)$, where the base case $\ell(x)=\ell(u)$ is given. If $\ell(x)>\ell(u)$, find $u\leq z\lessdot x$ and let $x^{\prime}\neq x$ be the unique element such that $z\lessdot x^{\prime}\lessdot y$. By Lemma 3.3 on $[z,y]$, $\operatorname{wt}(x,y)\in\mathrm{span}_{\mathbb{R}}\{\operatorname{wt}(z,x),\operatorname{wt}(z,x^{\prime})\}\subset V$. So the induction step is complete. ∎
Proposition 3.6.
Let $\alpha_{i}$ be a right descent of $v$. We have the following recursion:
$\operatorname{AD}(u,v)=\begin{cases}\operatorname{AD}(us_{i},vs_{i})&\text{ if}us_{i}<u,\\ \operatorname{AD}(u,vs_{i})+\mathbb{R}\cdot\operatorname{wt}(u,us_{i})&\text{ if }us_{i}>u.\end{cases}$
Proof.
We consider these two cases separately.
Case 1: $vs_{i}<v$ and $us_{i}<u$. We use induction on $\ell(v)-\ell(u)$, with the base case $v=u$ being clear. First, we show $\operatorname{AD}(u,v)\subseteq\operatorname{AD}(us_{i},vs_{i})$. By Lemma 3.5, it suffices to show that for all $v\gtrdot s_{\alpha}v\in[u,v]$, $\alpha\in\operatorname{AD}(us_{i},vs_{i})$. If $s_{\alpha}v\neq vs_{i}$, by Lemma 3.4, $s_{\alpha}v$ has a right descent at $s_{i}$. Now $v\geq s_{\alpha}v\geq u$ and all three have right descents at $s_{i}$, so $vs_{i}\geq s_{\alpha}vs_{i}\geq us_{i}$. Thus, $s_{\alpha}vs_{i}$ and $vs_{i}$ both lie in $[us_{i},vs_{i}]$, meaning that $\alpha=\operatorname{wt}(s_{\alpha}vs_{i},vs_{i})\in\operatorname{AD}(us_{i},vs_{i})$.
Now assume $vs_{i}=s_{\alpha}v$. Then $vs_{i}\neq u$ because of right descents. So $vs_{i}>u$. Choose any $z$ such that $vs_{i}\gtrdot z\geq u$. Suppose that the rank $2$ interval $[z,v]$ contains $v\gtrdot vs_{i},w\gtrdot z$. Let $\beta=\operatorname{wt}(z,vs_{i})\in\operatorname{AD}(u,vs_{i})\subseteq\operatorname{AD}(us_{i},vs_{i})$ and $\gamma=\operatorname{wt}(z,w)$. By Lemma 3.4, $w$ has a right descent at $s_{i}$. So $v>w>u$ and $vs_{i}>ws_{i}>us_{i}$. By induction hypothesis on $[u,w]$, $\gamma\in\operatorname{AD}(u,w)=\operatorname{AD}(us_{i},ws_{i})\subseteq\operatorname{AD}(us_{i},vs_{i})$. By Lemma 3.3, $\alpha\in\mathrm{span}_{\mathbb{R}}\{\beta,\gamma\}\subseteq\operatorname{AD}(us_{i},vs_{i})$ as desired.
Example 3.7 shows a specific scenario of this situation.
Dually, the $\supseteq$ direction is proved via the exact same argument by arguing from the cover relations incident to $us_{i}$ inside $[us_{i},vs_{i}]$.
Case 2: $vs_{i}<v$ and $us_{i}>u$. For the $\subseteq$ direction, take any $\alpha$ such that $s_{\alpha}u\gtrdot u$ and $s_{\alpha}u\leq v$. If $s_{\alpha}u=us_{i}$, then $\alpha=\operatorname{wt}(u,us_{i})$ belongs to the right hand side. If $s_{\alpha}u\nleq us_{i}$, by Lemma 3.4, $s_{\alpha}u$ does not have a right descent at $s_{i}$. The lifting property (Proposition 2.2) on $s_{\alpha}u<v$ gives $s_{\alpha}u\leq vs_{i}$ so $\alpha\in\operatorname{AD}(u,vs_{i})$ as desired. For the $\supseteq$ direction, we just need $\operatorname{wt}(u,us_{i})\in\operatorname{AD}(u,v)$ which is true as $us_{i}\leq v$ by the lifting property. ∎
Example 3.7.
Let $u=3412$, $v=1324$ and $i=2$. According to the first case of Proposition 3.6, $\operatorname{AD}(1324,3412)=\operatorname{AD}(1234,3142)$. Both intervals are shown in Figure 1. In particular, the rank sizes of $[1324,3412]$ are $1,4,4,1$ while the rank sizes of $[1234,3142]$ are $1,3,3,1$. The covers from the maximum in these two intervals are highlighted. The weights from $3412$ are $e_{1}-e_{3}$, $e_{2}-e_{3}$, $e_{1}-e_{4}$ and $e_{2}-e_{4}$ while the weights from $3142$ are $e_{1}-e_{3}$, $e_{2}-e_{3}$ and $e_{2}-e_{4}$ (from left to right). The same linear space is spanned by these two sets of weights.
12341234123412341324132413241324124312431243124321342134213421341342134213421342142314231423142321432143214321432314231423142314312431243124312414321432143214322413241324132413314231423142314232143214321432143412341234123412
Figure 1. The lower interval in the Bruhat order below $3412$
Proposition 3.8.
$\operatorname{AD}(u,v)$ has the following properties.
  1. (1)
    For any $u\leq w\lessdot v$, $\operatorname{AD}(u,v)=\operatorname{AD}(u,w)+\mathbb{R}\cdot\operatorname{wt}(w,v)$.
  2. (2)
    For any saturated chain $u=w^{(0)}\lessdot w^{(1)}\lessdot\cdots\lessdot w^{(\ell-1)}\lessdot w^{(\ell)}=v$,
    $\operatorname{AD}(u,v)=\mathrm{span}_{\mathbb{R}}\{\operatorname{wt}(w^{(i)},w^{(i+1)})\>|\>i=0,\ldots,\ell-1\}.$
  3. (3)
    For any $w\in[u,v]$, $\operatorname{AD}(u,v)$ is spanned by the labels of all cover relations incident to $w$ inside $[u,v]$.
Proof.
We apply two layers of induction: first on $\ell=\ell(v)-\ell(u)$ and second on $\ell(u)$. The base cases where $\ell=0,1$ are true by definition. Assume $\ell\geq 2$. Denote these statements as ($1_{\ell}$), ($2_{\ell}$) and ($3_{\ell}$) respectively.
($1_{\ell}$$\Longrightarrow$($2_{\ell}$). For any such saturated chain,
[…]
$=$ $\operatorname{AD}(u,w^{(\ell-1)})+\mathbb{R}\cdot\operatorname{wt}(w^{(\ell-1)},v)=\operatorname{AD}(u,v)$
by ($2_{\ell-1}$) and ($1_{\ell}$).
($2_{\ell}$$\Longrightarrow$($3_{\ell}$). By Lemma 3.5, the labels of the upper covers from $w$ inside $[u,v]$ span $\operatorname{AD}(w,v)$, which by (2), equals the span of a saturated chain $C$ from $w$ to $v$. Similarly, the labels of the lower covers from $w$ have the same span as a saturated chain $C^{\prime}$ from $u$ to $w$. Putting the two chains together and by ($2_{\ell}$), we obtain the desired equality.
Thus, we now focus on proving ($1_{\ell}$). Fix $w=s_{\alpha}v$. If $v$ has a right descent $\alpha_{i}$ such that $vs_{i}\neq w$, go to Case 1 and Case 2. Otherwise, $v$ must have a unique right descent $\alpha_{i}$ satisfying $vs_{i}=w=s_{\alpha}v$, and we go to Case 3.
Case 1: $vs_{i}\neq w$ and $us_{i}<u$. Note that we don’t require $vs_{i}\geq u$ here. By Lemma 3.4, both $vs_{i}$ and $s_{\alpha}v$ cover $s_{\alpha}vs_{i}$. Also, since $s_{\alpha}v>u$ and both have right descents at $\alpha_{i}$, $s_{\alpha}vs_{i}>us_{i}$. By Proposition 3.6 and induction hypothesis on $\ell(us_{i})<\ell(u)$,
$\operatorname{AD}(u,v)=\operatorname{AD}(us_{i},vs_{i})=\operatorname{AD}(us_{i},s_{\alpha}vs_{i})+\mathbb{R}\alpha=\operatorname{AD}(u,s_{\alpha}v)+\mathbb{R}\alpha.$
Case 2: $vs_{i}\neq w$ and $us_{i}>u$. In this case, by the lifting property (Proposition 2.2) on $u<v$, we have $u\leq vs_{i}$. Similarly as above, $s_{\alpha}v$ has a right descent at $\alpha_{i}$. By the lifting property on $s_{\alpha}v>u$, we have $s_{\alpha}vs_{i}\geq u$ and $s_{\alpha}v\geq us_{i}$. Let $\beta=\operatorname{wt}(u,us_{i})$. By induction hypothesis on $\ell(vs_{i})-\ell(u)<\ell(v)-\ell(u)$ and Proposition 3.6,
$\operatorname{AD}(u,v)=\operatorname{AD}(u,vs_{i})+\mathbb{R}\beta=\operatorname{AD}(u,s_{\alpha}vs_{i})+\mathbb{R}\alpha+\mathbb{R}\beta=\operatorname{AD}(u,s_{\alpha}v)+\mathbb{R}\alpha.$
Case 3: $vs_{i}=s_{\alpha}v$. We are in this case when $\alpha_{i}$ is the only right descent of $v$. Choose any $w\gtrdot z\geq u$ and let $w^{\prime}\neq w$ be the unique element such that $v\gtrdot w^{\prime}\gtrdot z$. Let $\beta=\operatorname{wt}(w^{\prime},v)$, $\gamma=\operatorname{wt}(z,w)$ and $\tau=\operatorname{wt}(z,w^{\prime})$. By Lemma 3.3, $\mathrm{span}_{\mathbb{R}}\{\alpha,\gamma\}=\mathrm{span}_{\mathbb{R}}\{\beta,\tau\}$. Since $\alpha_{i}$ is the only right descent of $v$, Case 1 and Case 2 can be applied to $u\leq w^{\prime}\lessdot v$, giving us
$\operatorname{AD}(u,v)=$ $\operatorname{AD}(u,w^{\prime})+\mathbb{R}\beta$
$=$ $\operatorname{AD}(u,z)+\mathbb{R}\tau+\mathbb{R}\beta$
$=$ $\operatorname{AD}(u,z)+\mathbb{R}\gamma+\mathbb{R}\alpha$
$=$ $\operatorname{AD}(u,w)+\mathbb{R}\alpha$
by induction hypothesis ($1_{\ell-1}$). ∎
Definition 3.9.
A Bruhat interval $[u,v]$ is toric if $\operatorname{ad}(u,v)=\ell(v)-\ell(u)$.
In light of Proposition 3.8(2), we always have $\operatorname{ad}(u,v)\leq\ell(v)-\ell(u).$
Remark.
Unfortunately, being a toric interval is not a combinatorial invariant: the Bruhat interval $[1324,4231]$ (an example given in Remark 4.15 [CS23b]) is isomorphic to the boolean lattice $B_{4}$, but is not toric; in the meantime, there exist many Bruhat intervals that are isomorphic to $B_{4}$ and are toric (for example $[12345,51234]$). The toric property is still very close to a combinatorial invariant as shown in the main result of [CS23b] that if $\ell(v)-\ell(u)\leq n$, then $[u,v]$ is toric if and only if $[u,v]$ is a lattice.
In fact, knowing which intervals are toric provides us with full control over $\operatorname{ad}(u,v)$.
Proposition 3.10.
$\operatorname{ad}(u,v)=\max_{w\in[u,v]}\ell(w)-\ell(u)$ where $[u,w]$ is toric.
The dual statement of Proposition 3.10:
$\operatorname{ad}(u,v)=\max_{w\in[u,v]}\ell(v)-\ell(w)\text{ where }[w,v]\text{ is toric},$
follows from the exact same argument. It is also clear from Proposition 3.10 that the following more symmetric-looking statement holds:
$\operatorname{ad}(u,v)=\max_{u\leq x\leq y\leq v}\ell(y)-\ell(x)\text{ where }[x,y]\text{ is toric}.$
Proof.
For any $w\in[u,v]$ such that $[u,w]$ is toric, $\ell(w)-\ell(u)=\operatorname{ad}(u,w)\leq\operatorname{ad}(u,v)$. To show that $\operatorname{ad}(u,v)\leq\max_{w\in[u,v]}\ell(w)-\ell(u)$ where $[u,w]$ is toric, we use induction on $\ell(v)-\ell(u)$. Note that when $[u,v]$ is toric, the statement holds, and that by Lemma 3.3, whenever $\ell(v)-\ell(u)\leq 2$, $[u,v]$ is toric. Now assume that $[u,v]$ is not toric. It suffices to show that there exists $v\gtrdot v^{\prime}\geq u$ such that $\operatorname{ad}(u,v)=\operatorname{ad}(u,v^{\prime})$, since in which case, by induction hypothesis,
$\operatorname{ad}(u,v)=\operatorname{ad}(u,v^{\prime})=\max_{\begin{subarray}{c}w\in[u,v^{\prime}]\\ [u,w]\text{ is toric}\end{subarray}}\ell(w)-\ell(u)\leq\max_{\begin{subarray}{c}w\in[u,v]\\ [u,w]\text{ is toric}\end{subarray}}\ell(w)-\ell(u)\leq\operatorname{ad}(u,v),$
as desired. We first establish a helpful result.
Claim 3.11.
Consider $x,y_{1},y_{2}\in[u,v]$ such that $v\gtrdot y\gtrdot x$ for $y\in\{y_{1},y_{2}\}$. If $\operatorname{ad}(u,v)-\operatorname{ad}(u,x)=1$, then there exists $y\in\{y_{1},y_{2}\}$ such that $\operatorname{AD}(u,v)=\operatorname{AD}(u,y)$.
Proof of claim.
Let the labels $\operatorname{wt}(y_{1},v)$, $\operatorname{wt}(y_{2},v)$, $\operatorname{wt}(x,y_{1})$, $\operatorname{wt}(x,y_{2})$ be $\alpha,\beta,\gamma,\tau$, respectively. Assume the opposite that $\operatorname{AD}(u,y)\subsetneq\operatorname{AD}(u,v)$ for $y\in\{y_{1},y_{2}\}$. Since $\operatorname{AD}(u,x)\subset\operatorname{AD}(u,y)$ and $\operatorname{ad}(u,v)-\operatorname{ad}(u,x)=1$, we must have that $\operatorname{AD}(u,x)=\operatorname{AD}(u,y)$ for $y\in\{y_{1},y_{2}\}$, meaning that $\gamma,\tau\in\operatorname{AD}(u,x)$. By Lemma 3.3, $\mathrm{span}_{\mathbb{R}}\{\gamma,\tau\}=\mathrm{span}_{\mathbb{R}}\{\alpha,\beta\}$, so $\alpha,\beta\in\operatorname{AD}(u,x)$. By Proposition 3.8(1), $\operatorname{AD}(u,v)=\operatorname{AD}(u,y_{1})+\mathbb{R}\alpha=\operatorname{AD}(u,x)$, contradicting $\operatorname{ad}(u,v)=\operatorname{ad}(u,x)=1$. ∎
Now, fix a chain $v=v_{0}\gtrdot v_{1}\gtrdot\cdots\gtrdot v_{k}=u$, and consider the sequence of integers $d_{i}=\operatorname{ad}(u,v_{i-1})-\operatorname{ad}(u,v_{i})\in\{0,1\}$ for $i=1,\ldots,k$. By Proposition 3.8(2), $\sum_{i=1}^{k}d_{i}=\operatorname{ad}(u,v)<k$ because $[u,v]$ is not toric. Thus, there must be some zeros among the $d_{i}^{\prime}s$. We are going to tweak this chain to move 00’s upwards.
If $d_{j}=1$ and $d_{j+1}=0$, let $v_{j}^{\prime}\neq v_{j}$ be the unique element in $[v_{j+1},v_{j-1}]$. As $\operatorname{ad}(u,v_{j-1})-\operatorname{ad}(u,v_{j+1})=1$, by Claim 3.11, there exists $y\in\{v_{j},v_{j}^{\prime}\}$ such that $\operatorname{AD}(u,y)=\operatorname{AD}(u,v_{j-1})$. Since $d_{j}=1$, $y\neq v_{j}$ and thus $y=v_{j}^{\prime}$. Replace $v_{j}$ by $v_{j}^{\prime}$ in the original chain, we see that $d_{j}^{\prime}=0$ and $d_{j+1}=1$. Continue this process of moving 00’s upwards, we will eventually end up with $v_{1}^{\prime}\lessdot v$ such that $d_{1}^{\prime}=0$, meaning that $\operatorname{ad}(u,v)=\operatorname{ad}(u,v_{1}^{\prime})$ as desired. ∎
Remark.
The material of this subsection can be generalized to any Coxeter groups with the exact same arguments, with a chosen Coxeter datum. We choose to stay within the realm of finite Weyl groups for consistency of notation throughout this paper.
3.2. The torus complexity via Deodhar decompositions
We investigate the torus action on Richardson varieties in detail. Our main technical tool is the Deodhar decomposition. We will follow the notation of [MR04].
Definition 3.12.
[MR04, Section 3] For an expression $\textbf{w}=q_{i_{1}}\cdots q_{i_{\ell}}$ where each $q_{i_{k}}\in S\cup\{1\}$, define $w_{(k)}=q_{i_{1}}\cdots q_{i_{k}}$ for $k=1,\ldots,\ell$. Set $w_{(0)}=1$. Define the following sets:
$J_{\mathbf{w}}^{+}:=$$\{k\in\{1,\ldots,\ell\}\>|\>w_{(k-1)}<w_{(k)}\},$
$J_{\mathbf{w}}^{\circ}:=$$\{k\in\{1,\ldots,\ell\}\>|\>w_{(k-1)}=w_{(k)}\},$
$J_{\mathbf{w}}^{-}:=$$\{k\in\{1,\ldots,\ell\}\>|\>w_{(k-1)}>w_{(k)}\}.$
For a reduced expression $\textbf{v}=s_{i_{1}}\cdots s_{i_{\ell}}$, a subexpression is obtained from v by replacing some of the factors with $1$. A subexpression $\mathbf{u}$ of $\mathbf{v}$ is called distinguished if $u_{(j)}\leq u_{(j-1)}s_{i_{j}}$ for all $j=1,\ldots,\ell$. In other words, if multiplication by $s_{i_{j}}$ (on the right) decreases the length of $u_{(j-1)}$, then we must have $u_{(j)}=u_{(j-1)}s_{i_{j}}$ in a distinguished subexpression.
For every simple root $\alpha_{i}\in\Delta$, there is a corresponding homomorphism $\varphi_{i}:\operatorname{SL}_{2}\rightarrow G$. Define the following elements
$x_{\alpha_{i}}(m)=\varphi_{i}\begin{pmatrix}1&m\\ 0&1\end{pmatrix},\quad y_{\alpha_{i}}(m)=\varphi_{i}\begin{pmatrix}1&0\\ m&1\end{pmatrix},\quad\dot{s}_{i}=\varphi_{i}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}.$
For the majority of this work, we abuse notation and refer to the coset representative of $w\in W=N_{G}(T)/T$ in $N_{G}(T)$ by $w$. In their work, Marsch and Rietsch fix specific coset representatives. To stay in line with their notation and constructions, in this self-contained section, we follow their notational conventions. For $s_{i}\in W$, we fix $\dot{s}_{i}$ as our coset representative of $s_{i}\in W$ in $N_{G}(T)$. For any $w\in W$ with $w=s_{i_{1}}\cdots s_{i_{\ell}}$, let $\dot{w}:=\dot{s}_{i_{1}}\cdots\dot{s}_{i_{\ell}}$ be our coset representative of $w$ in $N_{G}(T)$.
As introduced in Section 2.2, for each $\alpha\in\Phi$ there is an associated root subgroup $U_{\alpha}$ in $G$. The simple root subgroup $U_{\alpha_{i}}$ for $\alpha_{i}\in\Delta$ is the image of $x_{\alpha_{i}}$ in $G$, namely $\{x_{\alpha_{i}}(m)\mid m\in\mathbb{C}\}$. For $\alpha\in\Phi$, there exists $w\in W$ and $\alpha_{i}\in\Delta$ such that $\alpha=w(\alpha_{i})$. Then $U_{\alpha}=\dot{w}U_{\alpha_{i}}{\dot{w}}^{-1}$; it is the image of $x_{\alpha}(m):=\dot{w}x_{\alpha_{i}}(m){\dot{w}}^{-1}$ in $G$. In a similar way, define $y_{\alpha}(m):=\dot{w}y_{\alpha_{i}}(m){\dot{w}}^{-1}$. Then $y_{\alpha}(m)=x_{-\alpha}(m)$. It will be important later, that for any $t\in T$ and $m\in\mathbb{C}$,
(3) $tx_{\alpha}(m)=x_{\alpha}(\alpha(t)m)t.$
Definition 3.13.
[MR04, Definition 5.1] For a reduced word $\mathbf{v}=s_{i_{1}}\cdots s_{i_{\ell}}$ and a distinguished subword $\mathbf{u}$, define the following set
$G_{\mathbf{u},\mathbf{v}}=\left\{g=g_{1}g_{2}\cdots g_{\ell}\left|\begin{array}[]{ll}g_{k}=x_{\alpha_{i_{k}}}(m_{k}){\dot{s}_{i_{k}}}^{-1}&\text{ if $k\in J_{\mathbf{u}}^{-}$,}\\ g_{k}=y_{\alpha_{i_{k}}}(p_{k})&\text{ if $k\in J^{\circ}_{\mathbf{u}}$,}\\ g_{k}=\dot{s}_{i_{k}}&\text{ if $k\in J^{+}_{\mathbf{u}}$,}\end{array}\text{ for $p_{k}\in\mathbb{C}^{*},\,m_{k}\in\mathbb{C}$}\right.\right\}.$
We have $G_{\mathbf{u},\mathbf{v}}\simeq(\mathbb{C}^{*})^{J_{\mathbf{u}}^{\circ}}\times\mathbb{C}^{J_{\mathbf{u}}^{-}}$. Now, the Deodhar component is $D_{\mathbf{u},\mathbf{v}}:=G_{u,v}B/B\subset G/B$. We have the isomorphism $G_{u,v}\simeq D_{\mathbf{u},\mathbf{v}}$ via $g\mapsto gB/B$ ([MR04, Proposition 5.2 ]).
A distinguished subword $\mathbf{u}$ is called positive if $u_{(j-1)}\leq u_{(j)}$ for all $j$, i.e. $J_{\mathbf{u}}^{-}=\emptyset$. Given a fixed reduced word $\mathbf{v}$ of $v$ and an element $u\leq v$, there is a unique positive distinguished subword $\mathbf{u}_{+}$ of $u$, whose Deodhar component $D_{\mathbf{u}_{+},\mathbf{v}}$ has the same dimension as $\mathcal{R}_{u,v}$ by [MR04, Lemma 3.5].
Theorem 3.14.
[D85, Theorem 1.1] Given $u\leq v$ and fix a reduced word $\mathbf{v}$ of $v$. The Richardson cell has a decomposition $\mathcal{R}_{u,v}^{\circ}=\bigsqcup_{\mathbf{u}}D_{\mathbf{u},\mathbf{v}}$ into disjoint union, where the sum is over all distinguished subwords $\mathbf{u}$ of $\mathbf{v}$ that evaluate to $u$.
Example 3.15.
Let $v=321$ and $u=123$ in $S_{3}$. Pick a reduced word $\mathbf{v}=s_{1}s_{2}s_{1}$, and we have two distinguished subwords $s_{1}1s_{1}$ and $111$ for $u$. Among these two distinguished subwords, $111$ is positive. The components look like
$G_{s_{1}1s_{1},s_{1}s_{2}s_{1}}$ =$=\left\{\begin{pmatrix}&{-}1&\\ 1&&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}&1\end{pmatrix}\begin{pmatrix}{-}m_{3}&1&\\ 1&&\\ &&1\end{pmatrix}=\begin{pmatrix}{-}1&&\\ {-}m_{3}&1&\\ p_{2}&&1\end{pmatrix}\middle|\ p_{2}\in\mathbb{C}^{*},m_{3}\in\mathbb{C}\right\},$
$G_{111,s_{1}s_{2}s_{1}}$ =$=\left\{\begin{pmatrix}1&&\\ p_{1}&1&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}&1\end{pmatrix}\begin{pmatrix}1&&\\ p_{3}&1&\\ &&1\end{pmatrix}=\begin{pmatrix}1&&\\ p_{1}{+}p_{3}&1&\\ p_{2}p_{3}&p_{2}&1\end{pmatrix}\middle|\ p_{1},p_{2},p_{3}\in\mathbb{C}^{*}\right\}.$
Then we look at how the torus acts on specific points in the Deodhar components.
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\begin{pmatrix}t_{1}&&\\ &t_{2}&\\ &&t_{3}\end{pmatrix}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{pmatrix}&{-}1&\\ 1&&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}&1\end{pmatrix}\begin{pmatrix}{-}m_{3}&1&\\ 1&&\\ &&1\end{pmatrix}B/B$
$=$ $\begin{pmatrix}&{-}1&\\ 1&&\\ &&1\end{pmatrix}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\begin{pmatrix}t_{2}&&\\ &t_{1}&\\ &&t_{3}\end{pmatrix}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{pmatrix}1&&\\ &1&\\ &p_{2}&1\end{pmatrix}\begin{pmatrix}{-}m_{3}&1&\\ 1&&\\ &&1\end{pmatrix}B/B$
$=$ $\begin{pmatrix}&{-}1&\\ 1&&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}\frac{t_{3}}{t_{1}}&1\end{pmatrix}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\begin{pmatrix}t_{2}&&\\ &t_{1}&\\ &&t_{3}\end{pmatrix}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{pmatrix}{-}m_{3}&1&\\ 1&&\\ &&1\end{pmatrix}B/B$
$=$ $\begin{pmatrix}&{-}1&\\ 1&&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}\frac{t_{3}}{t_{1}}&1\end{pmatrix}\begin{pmatrix}{-}m_{3}\frac{t_{2}}{t_{1}}&1&\\ 1&&\\ &&1\end{pmatrix}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\begin{pmatrix}t_{1}&&\\ &t_{2}&\\ &&t_{3}\end{pmatrix}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B/B$
$=$ $\begin{pmatrix}&{-}1&\\ 1&&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}\frac{t_{3}}{t_{1}}&1\end{pmatrix}\begin{pmatrix}{-}m_{3}\frac{t_{2}}{t_{1}}&1&\\ 1&&\\ &&1\end{pmatrix}B/B,$
where the torus “disappears” from the right since $T\subset B$. Thus, we can say that $T$ acts on $D_{s_{1}1s_{1},s_{1}s_{2}s_{1}}$ by weights $e_{3}-e_{1}$ and $e_{2}-e_{1}$, which span a $2$-dimensional space in the root space. With a similar calculation,
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\begin{pmatrix}t_{1}&&\\ &t_{2}&\\ &&t_{3}\end{pmatrix}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{pmatrix}1&&\\ p_{1}&1&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}&1\end{pmatrix}\begin{pmatrix}1&&\\ p_{3}&1&\\ &&1\end{pmatrix}B/B$
$=$ $\begin{pmatrix}1&&\\ p_{1}\frac{t_{2}}{t_{1}}&1&\\ &&1\end{pmatrix}\begin{pmatrix}1&&\\ &1&\\ &p_{2}\frac{t_{3}}{t_{2}}&1\end{pmatrix}\begin{pmatrix}1&&\\ p_{3}\frac{t_{2}}{t_{1}}&1&\\ &&1\end{pmatrix}B/B,$
so $T$ acts on $D_{111,s_{1}s_{2}s_{1}}$ by weights $e_{2}-e_{1},e_{3}-e_{2},e_{2}-e_{1}$, whose span is $2$-dimensional.
Definition 3.16.
Fix a reduced word $\mathbf{v}$ of $v$ and a distinguished subword $\mathbf{u}$. For each $k\in J_{\mathbf{u}}^{\circ}$, define $\beta_{k}=\operatorname{wt}(u_{(k-1)},u_{(k-1)}s_{i_{k}})$; for each $k\in J_{\mathbf{u}}^{-}$, define $\beta_{k}=\operatorname{wt}(u_{(k)},u_{(k-1)})$ (cf. Definition 3.1). Note that in both cases, $s_{\beta_{k}}=u_{(k-1)}s_{i_{k}}u_{(k-1)}^{-1}$ so $\beta_{k}=\pm u_{(k-1)}(\alpha_{i_{k}})$. To be precise, when $k\in J_{\mathbf{u}}^{\circ}$, $u_{(k-1)}<u_{(k-1)}s_{i_{k}}$ and when $k\in J_{\mathbf{u}}^{-}$, $u_{(k-1)}>u_{(k-1)}s_{i_{k}}$, so we have that $\beta_{k}=u_{(k-1)}(\alpha_{i_{k}})$ when $k\in J_{\mathbf{u}}^{\circ}$ and $\beta_{k}=-u_{(k-1)}(\alpha_{i_{k}})$ when $k\in J_{\mathbf{u}}^{-}$. Also define
$U_{k}=\begin{cases}U^{*}_{-\beta_{k}}&\textrm{if }k\in J_{\mathbf{u}}^{\circ}\\ U_{-\beta_{k}}&\textrm{if }k\in J_{\mathbf{u}}^{-}\\ \end{cases}.$
Define $\operatorname{TD}(\mathbf{u},\mathbf{v})=\mathrm{span}_{\mathbb{R}}\{\beta_{k}\>|\>k\in J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}\}$ and $\operatorname{td}(\mathbf{u},\mathbf{v})=\dim\operatorname{TD}(\mathbf{u},\mathbf{v})$.
For any $x\in D_{\mathbf{u},\mathbf{v}}$, let $g\in G_{\mathbf{u},\mathbf{v}}$ be the standard form of $x$ if they lie in the same $B$-coset. For such a $g$, let
$g^{(k)}=\begin{cases}p_{k}&\textrm{if }k\in J_{\mathbf{u}}^{\circ}\\ m_{k}&\textrm{if }k\in J_{\mathbf{u}}^{-}\\ \end{cases}$
We say that $x$ is a general point in $D_{\mathbf{u},\mathbf{v}}$ if the corresponding $m_{k}$’s in its standard form are all nonzero (cf. Definition 3.13).
Lemma 3.17.
Fix $u\leq v$ and a reduced word $\mathbf{v}=s_{i_{1}}\cdots s_{i_{\ell}}$ of $v$. Let $\mathbf{u}=u_{1}\cdots u_{\ell}$ be a distinguished subword of $\mathbf{v}$ for $u$. Let $J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}=\{j_{1}<j_{2}<\cdots<j_{z}\}$ and $g\in G_{\mathbf{u},\mathbf{v}}$. Then
$g=\displaystyle\prod_{k=1}^{z}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u},\qquad D_{\mathbf{u},\mathbf{v}}=\displaystyle\prod_{k=1}^{z}U_{j_{k}}\dot{u}B.$
Proof.
The first equality directly implies the second, and hence we focus on proving the first via induction on $\ell$. When $\ell=1$, $u=u_{1}=u_{(1)}$. If $u_{1}=s_{i_{1}}$, then $J_{\mathbf{w}}^{+}=\{1\}$ and $g=\dot{s}_{i_{1}}$. Otherwise, if $u_{1}=1$, then $J_{\mathbf{w}}^{\circ}=\{1\}$ and $g=y_{\alpha_{i_{1}}}(g^{(1)})=x_{-\alpha_{i_{1}}}(g^{(1)})=x_{-\beta_{1}}(g^{(1)})$.
Assume $\ell\geq 2$. By the inductive hypothesis, applied to $\mathbf{v^{\prime}}=s_{i_{1}}\cdots s_{i_{\ell-1}}$ and $\mathbf{u^{\prime}}=u_{1}\cdots u_{\ell-1}$ a distinguished subword of $\mathbf{v^{\prime}}$ for $u_{(\ell-1)}$,
(4) $g_{1}\cdots g_{\ell-1}=\displaystyle\prod_{k=1}^{z^{\prime}}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}_{(\ell-1)},$
where $J_{\mathbf{u^{\prime}}}^{\circ}\cup J_{\mathbf{u^{\prime}}}^{-}=\{j_{1}<j_{2}<\cdots<j_{z^{\prime}}\}\subseteq J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}$ and $z^{\prime}$ either equals $z$ or $z-1$.
There are three possible cases to consider.
Case 1: $\ell\in J_{\mathbf{u}}^{+}$. In this case, $\dot{u}_{(\ell-1)}g_{\ell}=\dot{u}_{(\ell-1)}\dot{s}_{i_{\ell}}=\dot{u}_{(\ell)}=\dot{u}$. In addition, $J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}=J_{\mathbf{u^{\prime}}}^{\circ}\cup J_{\mathbf{u^{\prime}}}^{-}$ and $z=z^{\prime}$. The two preceding sentences, combined with (4), imply
$g=\displaystyle\prod_{k=1}^{z^{\prime}}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}_{(\ell-1)}g_{\ell}=\prod_{k=1}^{z}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}.$
Case 2: $\ell\in J_{\mathbf{u}}^{\circ}$. In this case, $\dot{u}_{(\ell-1)}g_{\ell}=\dot{u}_{(\ell-1)}y_{\alpha_{i_{\ell}}}(g^{(\ell)})=\dot{u}_{(\ell-1)}x_{-\alpha_{i_{\ell}}}(g^{(\ell)})$. Then
$\dot{u}_{(\ell-1)}x_{-\alpha_{i_{\ell}}}(g^{(\ell)})=\dot{u}_{(\ell-1)}x_{-\alpha_{i_{\ell}}}(g^{(\ell)}){\dot{u}_{(\ell-1)}}^{-1}\dot{u}_{(\ell-1)}=x_{-u_{(\ell-1)}(\alpha_{i_{\ell}})}(g^{(\ell)})\dot{u}_{(\ell-1)}=x_{-\beta_{\ell}}(g^{(\ell)})\dot{u},$
where the last equality follows from the case hypothesis and the definition of $\beta_{\ell}$. In addition, $J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}=J_{\mathbf{u^{\prime}}}^{\circ}\cup J_{\mathbf{u^{\prime}}}^{-}\cup\{\ell\}$ and $z=z^{\prime}+1$. Applying (4) yields
$g=\displaystyle\prod_{k=1}^{z^{\prime}}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}_{(\ell-1)}g_{\ell}=\prod_{k=1}^{z^{\prime}}x_{-\beta_{j_{k}}}(g^{(j_{k})})x_{-\beta_{\ell}}(g^{(\ell)})\dot{u}=\prod_{k=1}^{z}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}.$
Case 3: $\ell\in J_{\mathbf{u}}^{-}$. In this case, $\dot{u}_{(\ell-1)}g_{\ell}=\dot{u}_{(\ell-1)}x_{\alpha_{i_{\ell}}}(g^{(\ell)}){\dot{s}_{i_{\ell}}}^{-1}$. Then
$\dot{u}_{(\ell-1)}x_{\alpha_{i_{\ell}}}(g^{(\ell)}){\dot{s}_{i_{\ell}}}^{-1}$ $=\dot{u}_{(\ell-1)}x_{\alpha_{i_{\ell}}}(g^{(\ell)})s_{i_{\ell}}^{-1}(\dot{u}_{(\ell-1)}{\dot{s}_{i_{\ell}}}^{-1})^{-1}(\dot{u}_{(\ell-1)}{\dot{s}_{i_{\ell}}}^{-1})$
$=x_{u_{(\ell-1)}(\alpha_{i_{\ell}})}(g^{(\ell)})\dot{u}_{(\ell-1)}{\dot{s}_{i_{\ell}}}^{-1}$
$=x_{-\beta_{k}}(g^{(k)})\dot{u},$
where the last equality follows from $\beta_{k}\!=\!-u_{(k-1)}(\alpha_{i_{k}})$ when $k\!\in\!J_{\mathbf{u}}^{-}$ and the fact that $u=u_{(\ell)}=u_{(\ell-1)}s_{i_{\ell}}<u_{(\ell-1)}$ implies $\dot{u}_{(\ell)}=\dot{u}_{(\ell-1)}{\dot{s}_{i_{\ell}}}^{-1}$ [MR04, Proof of Proposition 5.2]. In addition, $J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}=J_{\mathbf{u^{\prime}}}^{\circ}\cup J_{\mathbf{u^{\prime}}}^{-}\cup\{\ell\}$ and $z=z^{\prime}+1$. Finally, by (4)
$g=\displaystyle\prod_{k=1}^{z^{\prime}}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}_{(\ell-1)}g_{\ell}=\prod_{k=1}^{z^{\prime}}x_{-\beta_{j_{k}}}(g^{(j_{k})})x_{-\beta_{\ell}}(g^{(\ell)})\dot{u}=\prod_{k=1}^{z}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}.$
As all three cases yield the desired formula for $g$ our proof is complete. ∎
Theorem 3.18.
Fix $u\leq v$ and fix a reduced word $\mathbf{v}$ of $v$. Let $\mathbf{u}$ be a distinguished subword of $\mathbf{v}$ for $u$. For a general point $x\in D_{\mathbf{u},\mathbf{v}}$ The span of the weights of the torus action on the orbit $Tx$ equals $\operatorname{TD}(\mathbf{u},\mathbf{v})$. And if $x\in D_{\mathbf{u},\mathbf{v}}$ is not general, then the torus weights on $Tx$ are contained in $\operatorname{TD}(\mathbf{u},\mathbf{v})$. Moreover, we have
$\operatorname{TD}(\mathbf{u},\mathbf{v})\subset\operatorname{TD}(\mathbf{u}_{+},\mathbf{v})=\operatorname{AD}(u,v).$
Proof.
The first part of the proof utilizes similar arguments as Theorem 4.7 of [CS23b], who analyzed the torus weights for positive distinguished subwords.
By Lemma 3.17, the $T$-action on $D_{\mathbf{u},\mathbf{v}}$ is given by the $T$-action on $\prod_{k=1}^{z}U_{j_{k}}\dot{u}B$ for $J_{\mathbf{u}}^{\circ}\cup J_{\mathbf{u}}^{-}=\{j_{1}<j_{2}<\cdots<j_{z}\}$. For $h\in T$ and $g$ the standard form of $x\in D_{\mathbf{u},\mathbf{v}}$,
$h\cdot g=h\cdot\prod_{k=1}^{z}x_{-\beta_{j_{k}}}(g^{(j_{k})})\dot{u}B=\prod_{k=1}^{z}x_{-\beta_{j_{k}}}(-\beta_{k}(h)g^{(j_{k})})\dot{u}B,$
[…]
It is now straightforward from the above arguments that if $x$ is not general, then the torus weights are contained in $\operatorname{TD}(\mathbf{u},\mathbf{v})$.
For the second part of the proof, we use induction on $\ell(v)$ to obtain a recurrence and relate it to recurrence of $\operatorname{AD}(u,v)$ (Proposition 3.6). Let $\mathbf{v}=s_{i_{1}}\cdots s_{i_{\ell}}$ and write $s_{i}=s_{i_{\ell}}$ for simplicity, as we will mainly be discussing this right descent of $v$. Write $\mathbf{u}=q_{i_{1}}\cdots q_{i_{\ell}}$ where $q_{i_{j}}\in\{s_{i_{j}},1\}$. Let $\mathbf{v}^{\prime}=s_{i_{1}}\cdots s_{i_{\ell-1}}$ and $\mathbf{u}^{\prime}=q_{i_{1}}\cdots q_{i_{\ell-1}}$. Prefixes of a distinguished subword are also distinguished so $\mathbf{u}^{\prime}$ is a distinguished subword of $\mathbf{v}^{\prime}$. Let $\mathbf{v}^{\prime}$ evaluate to $v^{\prime}=vs_{i}<v$ and let $\mathbf{u}^{\prime}$ evaluate to $u^{\prime}\in\{us_{i},u\}$.
Case 1: $us_{i}<u$. If $u^{\prime}=u$, then by the definition of a distinguished subword (Definition 3.12), we need to select $q_{i_{\ell}}=s_{i}$, contradicting $u=u^{\prime}q_{i_{\ell}}$. Thus, $u^{\prime}=us_{i}$, $q_{i_{\ell}}=s_{i}$ and $\ell\in J_{\mathbf{u}}^{+}$. The sequence of $\beta_{k}$’s (Definition 3.16) remains unchanged from $(\mathbf{u},\mathbf{v})$ to $(\mathbf{u}^{\prime},\mathbf{v}^{\prime})$ and $\mathbf{u}$ is positive if and only if $\mathbf{u}^{\prime}$ is positive. In addition, $\operatorname{AD}(u,v)=\operatorname{AD}(us_{i},vs_{i})$ (Proposition 3.6). Thus, the desired statement follows from $\operatorname{TD}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\subset\operatorname{TD}(\mathbf{u}_{+}^{\prime},\mathbf{v}^{\prime})=\operatorname{AD}(us_{i},vs_{i})$ by induction.
Case 2: $us_{i}>u$. Here, either $q_{i_{\ell}}=1$ or $q_{i_{\ell}}=s_{i}$ are possible for a distinguished subword $\mathbf{u}$. If we want $\mathbf{u}$ to be positive, then $q_{i_{\ell}}=1$. For both situations, $\beta_{\ell}=\operatorname{wt}(u,us_{i})$ and therefore $\operatorname{TD}(\mathbf{u},\mathbf{v})=\operatorname{TD}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})+\mathbb{R}\cdot\operatorname{wt}(u,us_{i})$. We further divide into subcases.
Subcase 2.1: $q_{i_{\ell}}=1$, $\ell\in J_{\mathbf{u}}^{\circ}$. In this case, $\mathbf{u}$ is positive if and only if $\mathbf{u}^{\prime}$ is positive. Thus, by induction hypothesis and the recurrence (Proposition 3.6),
$\operatorname{TD}(\mathbf{u}_{+},\mathbf{v}_{+})=\operatorname{TD}(\mathbf{u}_{+}^{\prime},\mathbf{v}_{+}^{\prime})+\mathbb{R}\beta_{\ell}=\operatorname{AD}(u,vs_{i})+\mathbb{R}\beta_{\ell}=\operatorname{AD}(u,v).$
Since $\operatorname{TD}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\subset\operatorname{TD}(\mathbf{u}_{+}^{\prime},\mathbf{v}_{+}^{\prime})$ by induction hypothesis, after adding $\mathbb{R}\beta_{\ell}$ on both sides, we have $\operatorname{TD}(\mathbf{u},\mathbf{v})\subset\operatorname{TD}(\mathbf{u}_{+},\mathbf{v}_{+})$ as desired.
Subcase 2.2: $q_{i_{\ell}}=s_{i}$, $\ell\in J_{\mathbf{u}}^{\circ}$. In this case, $\mathbf{u}$ is not positive and $\mathbf{u}^{\prime}$ is a distinguished subword for $u^{\prime}=us_{i}>u$. By induction hypothesis, $\operatorname{TD}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\subset\operatorname{AD}(us_{i},vs_{i})$. So
$\operatorname{TD}(\mathbf{u},\mathbf{v})=\operatorname{TD}(\mathbf{u}_{+}^{\prime},\mathbf{v}_{+}^{\prime})+\mathbb{R}\beta_{\ell}\subset\operatorname{AD}(us_{i},vs_{i})+\mathbb{R}\beta_{\ell}\subset\operatorname{AD}(u,vs_{i})+\mathbb{R}\beta_{\ell}=\operatorname{AD}(u,v).\qed$
Corollary 3.19.
For $u\leq v$, the Bruhat interval $[u,v]$ is toric if and only if the Richardson variety $\mathcal{R}_{u,v}$ is toric.
Proof.
By Theorem 3.18, the Richardson variety $\mathcal{R}_{u,v}$ is toric if $\dim\operatorname{TD}(\mathbf{u},\mathbf{v})=\ell(v)-\ell(u)$, for a fixed reduced word $\mathbf{v}$ of $v$ and some distinguished subword $\mathbf{u}$. This condition is equivalent to $\dim\operatorname{TD}(\mathbf{u}_{+},\mathbf{v})=\ell(v)-\ell(u)$ as $\operatorname{TD}(\mathbf{u}_{+},\mathbf{v})$ is maximal, which is equivalent to $\operatorname{ad}(u,v)=\ell(v)-\ell(u)$, i.e. $[u,v]$ is toric (c.f. Definition 3.9). ∎
We are now ready to prove the main theorem of this section, Theorem 1.2.
Proof of Theorem 1.2.
By Theorem 3.18, the maximum dimension of a $T$-orbit in the Richardson variety $\mathcal{R}_{u,v}$ is $\operatorname{ad}(u,v)$. By Definition 1.1, the $T$-complexity of $\mathcal{R}_{u,v}$ is
$c_{T}(\mathcal{R}_{u,v})=\ell(v)-\ell(u)-\operatorname{ad}(u,v)$
as desired. The second part of the theorem is Proposition 3.10. ∎
Proof of Corollary 1.3.
The subword property (Proposition 2.1) and Lemma 3.5 imply that $\operatorname{ad}(\mathrm{id},w)=\operatorname{supp}(w)$. The statement on the complexity of the Schubert variety $X_{w}$ now follows from Theorem 1.2. ∎
4. The Levi complexity of Schubert varieties
4.1. An equivariant isomorphism
In [GHY24], the authors of this work and A. Yong showed that if $w=w_{0}(I)d$ is length additive for an $I\subseteq\Delta$, then $X^{\circ}_{w}$ contains a distinguished subspace that is $T$-equivariantly isomorphic to $X^{\circ}_{d}$. The content in the first portion of this section is similar in broad form to our argument from [GHY24], but our results apply to any $w\in W$ and $I\subseteq\Delta$. This leads to a codimension preserving bijection between the $B_{L_{I}}$-orbits of $X_{w}^{\circ}$ and the $T$-orbits of $X_{\prescript{I}{}{w}}^{\circ}$.
Let $X$ be a $T$-variety with action denoted by $\boldsymbol{\cdot}$. For each $u\in W$ we define a $T$-action $\boldsymbol{\cdot}_{u}$ on $X$ by $t\boldsymbol{\cdot}_{u}x=utu^{-1}\boldsymbol{\cdot}x$ for all $x\in X$ and $t\in T$. The action $\boldsymbol{\cdot}_{u}$ does not depend on the coset representative chosen for $u$ and so is well-defined by [GHY24, Lemma 3.1]. Henceforth, $\boldsymbol{\cdot}$ will always denote the usual action of $T$ on $G/B$ by left multiplication.
For a positive integer $n$ let $\mathbb{A}^{n}$ be the affine n-space. We use the following well-known fact [B91, §14.12, §14.4] repeatedly in what follows. For all $v\in W$,
[…]
From now on, to simplify our notation let the left parabolic decomposition of $w$ with respect to $I\subseteq S$ be $w=ad$ where $a\in W_{I}$ and $d\in\prescript{I}{}{W}$. Recall the definition of $V_{d}$ in Subsection 2.2.
Definition 4.1.
The $I$-heart of the Schubert variety $X_{w}$, denoted $\operatorname{Heart}_{I}(X_{w})$, is the subvariety $V_{d}wB\subseteq U_{w}wB=X_{w}^{\circ}$.
Proposition 4.2.
$\operatorname{Heart}_{I}(X_{w})$ is $T$-stable for the action $\boldsymbol{\cdot}$. Additionally, $\operatorname{Heart}_{I}(X_{w})$ is $T$-equivariantly isomorphic to $X_{d}^{\circ}$ equipped with the $T$-action $\boldsymbol{\cdot}_{a^{-1}}$ under the map $\phi:\operatorname{Heart}_{I}(X_{w})\rightarrow X_{d}^{\circ}$ given by $hB\mapsto a^{-1}hB$.
Proof.
Let $t\in T$. Then
$t\boldsymbol{\cdot}\operatorname{Heart}_{I}(X_{w})=tV_{d}wB=t(t^{-1}V_{d}t)wB=V_{d}wB=\operatorname{Heart}_{I}(X_{w}),$
where the second equality follows from the fact that $V_{d}$ is normalized by $T$, and the third inequality from $wB$ being a $T$-fixed point.
The image of the map $\phi$ is $\phi(\operatorname{Heart}_{I}(X_{w}))=a^{-1}V_{d}wB=U_{d}dB=X_{d}^{\circ}$, where the second equality is the definition of $V_{d}$ and the third equality is (5). Thus $\phi$ is well-defined and surjective. As $\phi$ is left multiplication by $a^{-1}$, the map is also injective. Since (5) implies that $X_{d}^{\circ}$ is smooth, and hence normal, the preceding remarks along with Zariski’s main lemma imply that $\phi$ is an isomorphism of varieties. Now, let $t\in T$ and $hB\in\operatorname{Heart}_{I}(X_{w})$. We have
$\phi(t\boldsymbol{\cdot}hB)=a^{-1}thB=a^{-1}taa^{-1}hB=t\boldsymbol{\cdot}_{a^{-1}}\phi(hB),$
establishing the $T$-equivariance of $\phi$. ∎
Define $B_{L_{I}}:=L_{I}\cap B$ and $U_{L_{I}}$ to be the unipotent radical of $B_{L_{I}}$. Then $B_{L_{I}}=T\ltimes U_{L_{I}}$ is a Borel subgroup of $L_{I}$, and $U_{L_{I}}=B_{L_{I}}\cap U=\prod_{\alpha\in\Phi^{+}(I)}U_{\alpha}$ [B91, §14].
Set $U_{\bar{a}}=\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{\alpha}$. Then $U_{\bar{a}}$ is a closed subgroup of $U$ normalized by $T$ and
$B_{L_{I}}=T\ltimes U_{L_{I}}=T\ltimes\prod_{\alpha\in\phi^{+}(I)}U_{\alpha}=T\ltimes(U_{a}U_{\bar{a}}).$
For a group $H$ acting on a variety $X$, denote the isotropy subgroup of $H$ at $x\in X$ by $H_{x}$.
Lemma 4.3.
Let $x\in X_{w}^{\circ}$. Then $U_{\bar{a}}\leq B_{x}$.
Proof.
The big open cell $X_{w}^{\circ}$ is a $B$-orbit and so the isotropy subgroups $B_{y}$ are identical at every point in $y\in X_{w}^{\circ}$. Consider the isotropy subgroup at the point $wB\in X_{w}^{\circ}$. Then
$U_{\bar{a}}wB=ww^{-1}U_{\bar{a}}wB=w\prod_{\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{w^{-1}(\alpha)}B,$
where the last equality is [J03, Part II, 1.4(5)]. Let $\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)$. We claim $\alpha\notin{\mathcal{I}}(w)$. The claim would imply that $w^{-1}(\alpha)\in\Phi^{+}$, and continuing the above equation yields
$w\prod_{\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{w^{-1}(\alpha)}B=wB,$
proving that $U_{\bar{a}}\leq B_{wB}=B_{x}$.
We now show ${\mathcal{I}}(w)\cap\Phi^{+}(I)={\mathcal{I}}(a)$, which implies our claim and completes the proof. Now $a\in W_{I}$ implies ${\mathcal{I}}(a)\subseteq\Phi^{+}(I)$. This, combined with the fact that ${\mathcal{I}}(w)={\mathcal{I}}(a)\sqcup a({\mathcal{I}}(d))$ by (1), yields ${\mathcal{I}}(a)\subseteq{\mathcal{I}}(w)\cap\Phi^{+}(I)$. To show the other inclusion, we first show that $d^{-1}(\Phi^{+}(I))\subseteq\Phi^{+}$. The fact that $d\in\prescript{I}{}{W}$ implies ${\mathcal{D}}_{L}(w)\cap I=\emptyset$ which in turn indicates ${\mathcal{I}}(d)\cap I=\emptyset$. Hence $d^{-1}(\alpha_{i})\subseteq\Phi^{+}$ for all $\alpha_{i}\in I$. So $d^{-1}(\Phi^{+}(I))\subseteq\Phi^{+}$ as $I$ generates $\Phi^{+}(I)$.
Now let $\alpha\in{\mathcal{I}}(w)\cap\Phi^{+}(I)$. Suppose, for contradiction, that $a^{-1}(\alpha)\in\Phi^{+}$. This, combined with $a\in W_{I}$ and $\alpha\in\Phi^{+}(I)$, means that $a^{-1}(\alpha)\in\Phi^{+}(I)$. Then, $w^{-1}(\alpha)=d^{-1}(a^{-1}(\alpha))\in\Phi^{+}$ contradicting $\alpha\in{\mathcal{I}}(w)$. So $\alpha\in{\mathcal{I}}(a)$. ∎
As an immediate corollary we have
Corollary 4.4.
Let $x\in X_{w}^{\circ}$. Then $U_{\bar{a}}\leq(B_{L_{I}})_{x}$.
We continue to work towards an understanding of the form of $(B_{L_{I}})_{x}$ by first proving a technical lemma.
Lemma 4.5.
Let $x\in X_{w}^{\circ}\setminus\operatorname{Heart}_{I}(X_{w})$ and $h_{1},h_{2}\in\operatorname{Heart}_{I}(X_{w})$.
  1. (1)
    $uh_{1}\notin\operatorname{Heart}_{I}(X_{w})$ for all $u\in U_{a}$ with $u\neq e$.
  2. (2)
    $tx\notin\operatorname{Heart}_{I}(X_{w})$ for all $t\in T$.
  3. (3)
    Let $b=tu_{1}u_{2}\in B_{L_{I}}$ with $t\in T$, $u_{1}\in U_{a}$, and $u_{2}\in U_{\bar{a}}$. If $bh_{1}=h_{2}$, then $u_{1}=e$ and $th_{1}=h_{2}$.
Proof.
(1) Since $h_{1}\in\operatorname{Heart}_{I}(X_{w})$, $h_{1}=vwB$ for some $v\in V_{d}$. By Lemma 2.3, $uv\notin V_{d}$. Thus $uvwB\in X_{w}^{\circ}\setminus\operatorname{Heart}_{I}(X_{w})$ by (5).
(2) The $T$-stability of $\operatorname{Heart}_{I}(X_{w})$, proved in Proposition  4.2, implies $tx\notin\operatorname{Heart}_{I}(X_{w})$.
(3) By Lemma 4.3, $u_{2}h_{1}=h_{1}$. Suppose, for contradiction that $u_{1}\neq e$. Then (1) implies that $u_{1}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$. And thus, (2) implies $tu_{1}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$. This contradicts our assumption that $tu_{1}u_{2}h_{1}=h_{2}$. Thus $u_{1}=e$ and $th_{1}=bh_{1}=h_{2}$. ∎
Finally, we are able to describe the precise form of $(B_{L_{I}})_{x}$.
Lemma 4.6.
Let $x\in X_{w}^{\circ}$. Then $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$. And $(B_{L_{I}})_{x}=T_{h}\ltimes U_{\bar{a}}$ for any $h\in(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})$.
Proof.
Let $x\in X_{w}^{\circ}$. By (5), $x=uwB$ for $u\in U_{w}$. Thus, by Lemma 2.3, $x=u_{1}vwB$ for $u_{1}\in U_{a},v\in V_{d}$. Since $u_{1}^{-1}\in U_{a}\leq B_{L_{I}}$, we have $vwB\in B_{L_{I}}\cdot x$ and $vwB\in\operatorname{Heart}_{I}(X_{w})$ by definition. Thus $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$.
Let $h\in(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})$. It is immediate by Corollary 4.4 that $T_{h}\ltimes U_{\bar{a}}\leq(B_{L_{I}})_{h}$. On the other hand, let $b\in(B_{L_{I}})_{h}$ so that $bh=h$. We have $b=tu_{1}u_{2}$ with $t\in T$, $u_{1}\in U_{a}$, and $u_{2}\in U_{\bar{a}}$. Lemma 4.5(3) implies that $u_{1}=e$ and $th=h$. Hence $b\in T_{h}\ltimes U_{\bar{a}}$. We conclude that $(B_{L_{I}})_{h}=T_{h}\ltimes U_{\bar{a}}$. Our proof is complete since $(B_{L_{I}})_{x}=(B_{L_{I}})_{h}$. ∎
We now define our bijection between $B_{L_{I}}$-orbits in $X_{w}^{\circ}$ and $T$-orbits in $\operatorname{Heart}_{I}(X_{w})$.
Theorem 4.7.
The map $\beta:\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))\rightarrow\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\mapsto B_{L_{I}}x$ where $x$ is any point in $\Theta$ is a codimension-preserving bijection. That is,
$\dim(X_{w}^{\circ})-\dim(\beta(\Theta))=\dim(\operatorname{Heart}_{I}(X_{w}))-\dim(\Theta),$
for every $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$.
Proof.
We begin by noting that $T\leq B_{L_{I}}$ implies that this map is well-defined.
Let $\Theta_{1},\Theta_{2}\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ with $x\in\Theta_{1}$ and $y\in\Theta_{2}$. Suppose that $B_{L_{I}}\cdot x=B_{L_{I}}\cdot y$. This implies $b_{1}x=b_{2}y$ for some $b_{1},b_{2}\in B_{L_{I}}$. Thus $b_{2}^{-1}b_{1}x=y$. Lemma 4.5(3) implies there exists a $t\in T$ such that $tx=y$. Thus $\Theta_{1}=\Theta_{2}$ and $\beta$ is injective.
Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$. By Lemma 4.6, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$. Thus $h\in\Theta$ for some $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ and so $\beta(\Theta)=\Xi$. Thus $\beta$ is surjective. Continuing with this notation, we now consider the codimension of the orbits $\Xi$ and $\Theta$ in $X_{w}^{\circ}$ and $\operatorname{Heart}_{I}(X_{w})$, respectively. One has, by [B09, Proposition 1.11], that for any algebraic group $H$ and $H$-variety $X$, the orbit $H\cdot x$, $x\in X$, is a subvariety of $X$ of dimension $\dim H-\dim H_{x}$. Thus
$\dim(\Xi)$ $=\dim(B_{L_{I}})-\dim(B_{L_{I}})_{x}$
$=\dim(B_{L_{I}})-\dim(T_{h}\ltimes U_{\bar{a}})$
$=(\ell(w_{0}(I))+\dim(T))-((\dim(T)-\dim(\Theta))+(\ell(w_{0}(I))-\ell(a)))$
$=\ell(a)+\dim(\Theta),$
where the second equality is Lemma 4.6, and the third equality follows from the definition of $B_{L_{I}}$ and $U_{\bar{a}}$ as well as [B09, Proposition 1.11] applied to the $T$-orbit $\Theta$. Hence
$\dim(X_{w}^{\circ})-\dim(\Xi)$ $=(\ell(a)+\ell(d))-(\ell(a)+\dim(\Theta))$
$=\ell(d)-\dim(\Theta)$
$=\dim(\operatorname{Heart}_{I}(X_{w}))-\dim(\Theta),$
where the first equality is the length additivity of $ad$, while the final equality follows from Proposition 4.2. ∎
We are now able to prove Theorem 1.4.
Proof of Theorem 1.4 Set $a={{}_{I}w}$ and $d=\prescript{I}{}{w}$ to match the notation we have been using throughout this section. Then [GHY24, Lemma 3.2] tells us that the set of $T$-orbits in $X^{\circ}_{d}$ for the action $\boldsymbol{\cdot}_{a^{-1}}$ is identical to the set of $T$-orbits in $X^{\circ}_{d}$ for the action $\boldsymbol{\cdot}$. Hence, Proposition 4.2 implies that the map $\mathfrak{T}:\mathcal{O}_{T}(X^{\circ}_{d})\rightarrow\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ given by $\Theta\mapsto Tax$, where $x$ is any point in $\Theta$, is a codimension preserving bijection. The map $\mathfrak{O}$ is precisely the composition of two codimension-preserving bijections, namely $\mathfrak{O}=\beta\circ\mathfrak{T}$. Hence $\mathfrak{O}$ is itself a codimension-preserving bijection. ∎
Remark.
For the remainder of this paper, we abuse notation slightly by writing $C_{B_{L_{I}}}(X^{\circ}_{w})$ and $C_{B_{L_{I}}}(X_{w})$ for the minimum codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ and $X_{w}$, respectively. This allows for discussing these orbit codimensions even when $L_{I}$ does not act on $X^{\circ}_{w}$ or $X_{w}$.
Corollary 4.8.
The minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ equals the torus complexity of $X^{\circ}_{d}$ for the usual torus action. That is, $c_{B_{L_{I}}}(X^{\circ}_{w})=c_{T}(X^{\circ}_{d})$.
4.2. Orbits in the Schubert Variety
In this section let $w\in W$ and $I\subseteq\Delta$. For any $u\in W$, let $u={{}_{I}u}\prescript{I}{}{u}$ be the left parabolic decomposition of $u$ with respect to $I$. The goal of this section is to use Theorem 4.7 to prove Corollary 1.6 and Theorem 1.5.
Proof of Corollary 1.6. The Bruhat decomposition tells us that the Schubert variety $X_{w}$ is the disjoint union of the Schubert cells for $u\leq w$,
$X_{w}=\displaystyle\bigsqcup_{u\leq w}X_{u}^{\circ}.$
and hence
(6) $\mathcal{O}_{B_{L_{I}}}(X_{w})=\bigsqcup_{u\leq w}\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u}).$
Theorem 4.7 implies there is a bijection between $\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})$ and $\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})\times\{u\}$ for all $u\leq w$, which completes our proof. ∎
Remark.
In Corollary 1.6 we decorate each of the sets $\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})$ by taking it’s product with $\{u\}$. This is due to the fact that for $u,v\leq w$ with $u\neq v$ it is possible that $\prescript{I}{}{u}=\prescript{I}{}{v}$. And we desire the union in the statement of Corollary 1.6 to be disjoint.
Our next order of business is to apply Corollary 4.8 towards a formula for the minimal codimension of a $B_{L_{I}}$-orbit in any Schubert variety $X_{w}$.
Proposition 4.9.
For $w\in W$, the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$ is given by
$c_{B_{L_{I}}}(X_{w})=\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right).$
Proof.
For a $u\leq w$, the minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{u}$ is $\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})$ by Corollary 4.8 and Corollary 1.3. The codimension of $X^{\circ}_{u}$ in $X_{w}$ is $\ell(w)-\ell(u)$. These two preceding sentences combine with (6) to imply
$c_{B_{L_{I}}}(X_{w})=$ $\min_{u\leq w}\left(\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})+\ell(w)-\ell(u)\right)$
$=$ $\ell(w)+\min_{u\leq w}\left(-\ell(_{I}u)-\operatorname{supp}(\prescript{I}{}{u})\right)$
$=$ $\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right),$
where the second inequality is the length-additivity of the left parabolic decomposition. ∎
The following example illustrates that this value is not, in general, well-behaved.
Example 4.10.
Let $w=56781234\in S_{8}$ written in one-line notation. Then $w$ is already in $\prescript{I}{}{W}$ since $\mathcal{D}_{L}(w)\cap I$ is the empty set. Therefore, $\ell(_{I}w)+\operatorname{supp}(\prescript{I}{}{w})=0+7=7$. Now if we apply $s_{4}$ to $w$ and get $u=56718234$, then $\prescript{I}{}{u}=25681347$, and ${}_{I}u$ has length $4$. Therefore, $\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})=4+7=11$. This is not the optimal choice for $u$: if we take $u=53218764\leq w$, then $\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})=6+7=13$.
However, the formula in Proposition 4.9 can be simplified in restricted cases.
Lemma 4.11.
Let $w\in W$ such that $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$, i.e. $w$ admits a left Billey-Postnikov decomposition with respect to $I$ (see [RS17, Proposition 4.2]). Then $w$ simultaneously maximizes both $\ell(_{I}u)$ and $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$. Thus
$c_{B_{L_{I}}}(X_{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
Proof.
For any $w$, if $u\leq w$, then $\prescript{I}{}{u}\leq\prescript{I}{}{w}$. Thus $w$ always maximizes $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$. Hence, if $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$, then $w$ simultaneously maximizes both $\ell(_{I}u)$ and $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$. Applying this to Proposition 4.9 yields
$c_{B_{L_{I}}}(X_{w})=$ $\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right)$
$=$ $\ell(w)-\ell(_{I}w)-\operatorname{supp}(\prescript{I}{}{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
The hypothesis of Lemma 4.11 is satisfied when $L_{I}$ acts $X_{w}$. This yields one of our main results, Theorem 1.5: a simple, type independent formula for the Levi complexity of a Schubert variety. We are now ready to prove Theorem 1.5.
Proof of Theorem 1.5.
The equivalence of $L_{I}$ acting on $X_{w}$ and $I\subseteq\mathcal{D}_{L}(w)$ is [BL00, Lemma 8.2.3]. If $I\subseteq{\mathcal{D}}_{L}(w)$, then ${\mathcal{I}}(w)$ contains all the positive roots in the root subsystem generated by $I$. Thus $w=w_{0}(I)d$ is a length-additive expression for some $d\in W$ by [BB05, Proposition 3.1.3]. That is, ${}_{I}w=w_{0}(I)$ in the left parabolic decomposition of $w$. Thus, it must be the case then that $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$. Hence Lemma 4.11 implies the theorem statement. ∎
4.3. Complexity in the partial flag variety
For $J\subset\Delta$, $G/P_{J}$ is called a partial flag variety. For $w\in W^{J}$, the $B$-orbit of $wP_{J}$ is referred to as the Schubert cell $X^{J,\circ}_{w}$; the closure of $X^{J,\circ}_{w}$ in $G/P_{J}$ with respect to the Zariski topology is the Schubert variety $X^{J}_{w}$. When $J=\emptyset$, then $P_{J}=B$ and $G/B$ is the full flag variety. In this case, for $w\in W^{\emptyset}=W$ we shall continue to write $X_{w}^{\circ}$ and $X_{w}$ for the Schubert cell and variety in $G/B$, respectively. The goal of this section is to extend, where possible, the Levi-Borel complexity results of Section 4.2 to Schubert varieties in the partial flag variety.
We recall the relationship between Schubert varieties in the partial flag variety and Schubert varieties in the full flag variety. For any $w\in W^{J}$, the canonical projection $\pi_{B,P_{J}}:G/B\rightarrow G/P_{J}$ is $G$-equivariant and restricts to a birational morphism $\pi_{B,P_{J}}|_{X_{w}}:X_{w}\rightarrow X^{P_{J}}_{w}$. The inverse image is $\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w})=X_{ww_{0}(J)}$.
Lemma 4.12.
Let $w\in W^{J}$. Then ${\rm stab}_{G}(X^{P_{J}}_{w})=P_{D_{L}(ww_{0}(J))}$.
Proof.
We have ${\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w}))\!=\!{\rm stab}_{G}(X_{ww_{0}(J)})\!=\!P_{D_{L}(ww_{0}(J))}$ by [BL00, Lemma 8.2.3]. Thus our result will follow if we show
${\rm stab}_{G}(X^{P_{J}}_{w})={\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w})).$
We first show ${\rm stab}_{G}(X^{P_{J}}_{w})\subseteq{\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w}))$. Let $g\in{\rm stab}_{G}(X^{P_{J}}_{w})$ and $x\in\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w})$. Then $\pi_{B,P_{J}}(x)\in X^{P_{J}}_{w}$, and thus $g\in{\rm stab}_{G}(X^{P_{J}}_{w})$ implies $g\pi_{B,P_{J}}(x)\in X^{P_{J}}_{w}$. The $G$-equivariance of $\pi_{B,P_{J}}$ yields $\pi_{B,P_{J}}(gx)\in X^{P_{J}}_{w}$, which implies $gx\in\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w})$. We conclude $g\in{\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w}))$.
Now we verify ${\rm stab}_{G}(X^{P_{J}}_{w})\supseteq{\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w}))$. Let $g\in{\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w}))$, $x\in X^{P_{J}}_{w}$, and $z\in\pi_{B,P_{J}}^{-1}(x)$. Then $gx=g\pi_{B,P_{J}}(z)=\pi_{B,P_{J}}(gz)$ by the $G$-equivariance of the map. Now since $g\in{\rm stab}_{G}(\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w}))$, we have $gz\!\in\!\pi_{B,P_{J}}^{-1}(X^{P_{J}}_{w})$. As desired, we have $gx=\pi_{B,P_{J}}(gz)\in X^{P_{J}}_{w}$ and thus $g\in{\rm stab}_{G}(X^{P_{J}}_{w})$. ∎
Theorem 4.13.
Let $w\in W^{J}$ and $I\subseteq D_{L}(ww_{0}(J))$. Then $L_{I}$ acts on $X^{P_{J}}_{w}$. If $L_{I}$ also acts on $X_{w}$ (equivalently if $I\subseteq P_{D_{L}(w)}$), then $c_{L_{I}}(X^{P_{J}}_{w})=c_{L_{I}}(X_{w})$.
Proof.
The Levi subgroup $L_{I}$ acts on $X^{P_{J}}_{w}$ by Lemma 4.12. If $I\subseteq P_{D_{L}(w)}$, then $L_{I}$ acts on $X_{w}$. Thus, since $\pi_{B,P_{J}}:G/B\rightarrow G/P_{J}$ is $G$-equivariant, we have that $\pi_{B,P_{J}}|_{X_{w}}:X_{w}\rightarrow X^{P_{J}}_{w}$ is an $L_{I}$-equivariant birational morphism. And $L_{I}$-complexity is invariant under $L_{I}$-equivariant birational maps[P10, §2]. Thus $c_{L_{I}}(X^{P_{J}}_{w})=c_{L_{I}}(X_{w})$. ∎
We highlight that this does not reduce the Levi complexity of Schubert varieties in the partial flag variety to that of the full flag variety. This reduction only works if $L_{I}$ acts on both $X^{P_{J}}_{w}$ and $X_{w}$, and importantly ${\rm stab}_{G}(X_{w})$ is only a subgroup of ${\rm stab}_{G}(X^{P_{J}}_{w})$. However, this does resolve the case of $L_{I}=T$, since $T$ acts on any Schubert variety in any partial flag variety.
Corollary 4.14.
Let $w\in W^{J}$. Then $c_{T}(X^{P_{J}}_{w})=c_{T}(X_{w})$. In particular, $X^{P_{J}}_{w}$ is a toric variety if and only if each $s_{j}\in S$ appears at most once in any reduced expression of $w$, or equivalently, $\ell(w)=\operatorname{supp}(w)$.
Proof.
Let $I=\emptyset$. Then $L_{I}=T$ and $I\subseteq P_{D_{L}(ww_{0}(J))}$ and $I\subseteq P_{D_{L}(w)}$. Thus, by Theorem 4.13, $c_{T}(X^{P_{J}}_{w})=c_{T}(X_{w})$. ∎
Acknowledgements
We thank Alexander Yong for many helpful conversations throughout the development of this paper.

### v2

@@
 In the process, we give an explicit description of the torus action on any Deodhar component as well as describe the root subgroups that comprise the component.
-Secondly, we exhibit a codimension preserving bijection between Levi-Borel subgroup (a Borel subgroup of a Levi subgroup in a reductive algebraic group) orbits in the big open cell of a Schubert variety and torus orbits in the big open cell of a distinguished Schubert subvariety.
-This bijection has many applications: a type-uniform formula for the Levi-Borel complexity of the usual Levi-Borel subgroup action on a Schubert variety as well as an indexing set for Levi-Borel subgroup orbits in a Schubert variety.
+Secondly, when a Levi subgroup in a reductive algebraic group acts on a Schubert variety, we exhibit a codimension preserving bijection between the Levi-Borel subgroup (a Borel subgroup in the Levi subgroup) orbits in the big open cell of that Schubert variety and torus orbits in the big open cell of a distinguished Schubert subvariety.
+This bijection has many applications including a type-uniform formula for the Levi-Borel complexity of the usual Levi-Borel subgroup action on a Schubert variety.
 We conclude by extending the Levi-Borel complexity results to a large class of Schubert varieties in the partial flag variety.
@@
 The homogeneous space $G/B$ is a smooth projective variety known as the full flag variety.
-The study of the flag variety first arose out of need to formalize and justify the enumerative geometry of H.
+The study of the flag variety first arose out of the need to formalize and justify the enumerative geometry of H.
 Schubert as laid out in Hilbert’s 15th problem.
@@
 Our work uncovers a deep structural relationship between the Levi-Borel orbits in the big cell of a Schubert variety and the $T$-orbits in the big cell of a distinguished Schubert subvariety.
-Specifically, we exhibit a codimension-preserving bijection between these sets of orbits.
+Specifically, when a Levi subgroup acts on a Schubert variety we exhibit a codimension-preserving bijection between these sets of orbits.
 Our choice of $T$ and $B$ determine the root system $\Phi$ and a set of simple roots $\Delta=\{\alpha_{1},\ldots,\alpha_{r}\}$, respectively.
@@
 Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$.
-The map $\mathfrak{O}\!:\!\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{w}})\!\rightarrow\!\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\!\mapsto\!B_{L_{I}}\prescript{I}{}{w}x$, where $x$ is any point in $\Theta$, is a codimension-preserving bijection.
+The map $\mathfrak{O}\!:\!\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{w}})\!\rightarrow\!\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\!\mapsto\!B_{L_{I}}\prescript{I}{}{w}x$, where $x$ is any point in $\Theta$, is a surjection.
+If $L_{I}$ acts on $X_{w}$, then $\mathfrak{O}$ is a codimension preserving bijection.
 That is,
@@
 Theorem 1.4 allows us to intertwine the study of Levi-Borel orbits with the vast literature on torus orbits and varieties, including our own Corollary 1.3, and has myriad applications which we now detail.
-Our first application is Proposition 4.9 which provides a formula for the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$.
-In general, the formula given in Proposition 4.9 is non-trivial to compute.
-However, in the case where the Levi subgroup $L_{I}$ acts on the Schubert variety $X_{w}$ the formula simplifies to a closed form.
+Our first application is Proposition 4.9 which provides a lower bound on the codimension of a $B_{L_{I}}$-orbit in $X_{w}$.
+In the case where the Levi subgroup $L_{I}$ acts on the Schubert variety $X_{w}$ this leads to a closed formula for the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$.
 While $B_{L_{I}}$ acts on any $X_{w}$, since $B_{L_{I}}\subseteq B$, the same is not true for $L_{I}$.
@@
 The interested reader may see [HY22, Theorem 2.4] for the details.
-As a second application of Theorem 1.4, we are able to describe the orbits of Levi-Borel subgroups in the Schubert variety itself.
-Corollary 1.6.
-Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$.
-There is a bijection between $\mathcal{O}_{B_{L_{I}}}(X_{w})$ and
-$\bigsqcup_{u\leq w}\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})\times\{u\}$
-We highlight that in the $I=\Delta$ case, $B_{L_{I}}=B$ and $\prescript{I}{}{u}=\mathrm{id}$ so each $\mathcal{O}_{T}(\prescript{I}{}{u})$ is a singleton set.
-Hence, in this case Corollary 1.6 recovers the Bruhat decomposition, namely that the Borel subgroup orbits in a Schubert variety are in bijection with the set of all $u\leq w$.
-In this way, Theorem 1.4 and Corollary 1.6 can be viewed as the first steps towards a framework that aims to generalize the study of Borel subgroup orbits in the full flag variety to the study of Levi-Borel subgroup orbits in Schubert varieties.
-Our next step, with work in progress, is understanding the inclusion order on these orbit closures.
-The combinatorial indexing set developed in this paper is an important cornerstone of this work.
+Our next step, with work in progress, is to use Theorem 1.4 to index the orbits of Levi-Borel subgroups in the Schubert variety itself and understand the inclusion order on their orbit closures.
 Throughout the years, the study of such Bruhat orders has led to much fruitful research.
@@
 In Section 3, we fully develop the notion of algebraic dimensions and establish Theorem 1.2.
-In Section 4, we prove Theorem 1.4, Theorem 1.5, and Corollary 1.6 with one of our main tools being an equivariant isomorphism that allows us to use results in Section 3.
+In Section 4, we prove Theorem 1.4 and Theorem 1.5 with one of our main tools being an equivariant isomorphism that allows us to use results in Section 3.
 Section 4 concludes with an extension of the Borel-Levi subgroup complexity results to Schubert varieties in partial flag varieties.
 2. Preliminaries
-We first recall some preliminaries and set notation which will be useful throughout.
 2.1. Combinatorics of the Weyl group
@@
 respectively. And its left inversion set and the right inversion set are
-${\mathcal{I}}_{L}(w)=\{\alpha\in\Phi^{+}\>|\>w(\alpha)\in\Phi^{-}\},\quad{\mathcal{I}}_{R}(w)=\{\alpha\in\Phi^{+}\>|\>w^{-1}(\alpha)\in\Phi^{-}\},$
+${\mathcal{I}}_{L}(w)=\{\alpha\in\Phi^{+}\>|\>w^{-1}(\alpha)\in\Phi^{-}\},\quad{\mathcal{I}}_{R}(w)=\{\alpha\in\Phi^{+}\>|\>w(\alpha)\in\Phi^{-}\},$
 respectively, with $|\mathcal{I}_{L}(w)|=|\mathcal{I}_{R}(w)|=\ell(w)$. There are equivalent ways to write down the inversion sets. For example,
@@
 3. The torus complexity of Richardson varieties
-In this section, we establish a formula for the complexity of the usual torus action on any Richardson variety $\mathcal{R}_{u,v}$ where $u\leq v$ in the Bruhat order.
+In this section, we establish a formula for the complexity of the usual torus $T$ action on any Richardson variety $\mathcal{R}_{u,v}$ where $u\leq v$ in the Bruhat order.
 Definition 3.1.
@@
 Then $B_{L_{I}}=T\ltimes U_{L_{I}}$ is a Borel subgroup of $L_{I}$, and $U_{L_{I}}=B_{L_{I}}\cap U=\prod_{\alpha\in\Phi^{+}(I)}U_{\alpha}$ [B91, §14].
-Set $U_{\bar{a}}=\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{\alpha}$. Then $U_{\bar{a}}$ is a closed subgroup of $U$ normalized by $T$ and
-$B_{L_{I}}=T\ltimes U_{L_{I}}=T\ltimes\prod_{\alpha\in\phi^{+}(I)}U_{\alpha}=T\ltimes(U_{a}U_{\bar{a}}).$
-For a group $H$ acting on a variety $X$, denote the isotropy subgroup of $H$ at $x\in X$ by $H_{x}$.
+Set $U_{\bar{a}}=\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{\alpha}$.
+If $\phi^{+}(I)\setminus{\mathcal{I}}(a)=\emptyset$, then we define $U_{\bar{a}}$ to be the trivial subgroup of $G$.
+Then $U_{\bar{a}}$ is a closed subgroup of $U$ normalized by $T$ and
+$B_{L_{I}}=T\ltimes U_{L_{I}}=T\ltimes\prod_{\alpha\in\phi^{+}(I)}U_{\alpha}=T\ltimes(U_{\bar{a}}U_{a}).$
 Lemma 4.3.
-Let $x\in X_{w}^{\circ}$. Then $U_{\bar{a}}\leq B_{x}$.
+The followings are equivalent:
+  1. (1)
+    $U_{\bar{a}}$ is the trivial subgroup of $G$.
+  2. (2)
+    $\phi^{+}(I)\setminus{\mathcal{I}}(a)=\emptyset$.
+  3. (3)
+    $L_{I}$ acts on $X_{w}$.
+  4. (4)
+    $I\subseteq\mathcal{D}_{L}(w)$.
+  5. (5)
+    $a=w_{0}(I)$.
 Proof.
-The big open cell $X_{w}^{\circ}$ is a $B$-orbit and so the isotropy subgroups $B_{y}$ are identical at every point in $y\in X_{w}^{\circ}$.
-Consider the isotropy subgroup at the point $wB\in X_{w}^{\circ}$.
-Then
-$U_{\bar{a}}wB=ww^{-1}U_{\bar{a}}wB=w\prod_{\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{w^{-1}(\alpha)}B,$
-where the last equality is [J03, Part II, 1.4(5)].
-Let $\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)$.
-We claim $\alpha\notin{\mathcal{I}}(w)$.
-The claim would imply that $w^{-1}(\alpha)\in\Phi^{+}$, and continuing the above equation yields
-$w\prod_{\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{w^{-1}(\alpha)}B=wB,$
-proving that $U_{\bar{a}}\leq B_{wB}=B_{x}$.
-We now show ${\mathcal{I}}(w)\cap\Phi^{+}(I)={\mathcal{I}}(a)$, which implies our claim and completes the proof.
-Now $a\in W_{I}$ implies ${\mathcal{I}}(a)\subseteq\Phi^{+}(I)$.
-This, combined with the fact that ${\mathcal{I}}(w)={\mathcal{I}}(a)\sqcup a({\mathcal{I}}(d))$ by (1), yields ${\mathcal{I}}(a)\subseteq{\mathcal{I}}(w)\cap\Phi^{+}(I)$.
-To show the other inclusion, we first show that $d^{-1}(\Phi^{+}(I))\subseteq\Phi^{+}$.
-The fact that $d\in\prescript{I}{}{W}$ implies ${\mathcal{D}}_{L}(w)\cap I=\emptyset$ which in turn indicates ${\mathcal{I}}(d)\cap I=\emptyset$.
-Hence $d^{-1}(\alpha_{i})\subseteq\Phi^{+}$ for all $\alpha_{i}\in I$.
-So $d^{-1}(\Phi^{+}(I))\subseteq\Phi^{+}$ as $I$ generates $\Phi^{+}(I)$.
-Now let $\alpha\in{\mathcal{I}}(w)\cap\Phi^{+}(I)$.
-Suppose, for contradiction, that $a^{-1}(\alpha)\in\Phi^{+}$.
-This, combined with $a\in W_{I}$ and $\alpha\in\Phi^{+}(I)$, means that $a^{-1}(\alpha)\in\Phi^{+}(I)$.
-Then, $w^{-1}(\alpha)=d^{-1}(a^{-1}(\alpha))\in\Phi^{+}$ contradicting $\alpha\in{\mathcal{I}}(w)$.
-So $\alpha\in{\mathcal{I}}(a)$.
-∎
-As an immediate corollary we have
-Corollary 4.4.
-Let $x\in X_{w}^{\circ}$. Then $U_{\bar{a}}\leq(B_{L_{I}})_{x}$.
-We continue to work towards an understanding of the form of $(B_{L_{I}})_{x}$ by first proving a technical lemma.
+One direction of the equivalence between (1) and (2) is by construction and the other follows from the definition of root subgroups.
+The equivalence of (3) and (4) is [BL00, Lemma 8.2.3].
+If $I\subseteq{\mathcal{D}}_{L}(w)$, then ${\mathcal{I}}(w)$ contains all the positive roots in the root subsystem generated by $I$.
+Thus $w=w_{0}(I)d$ is a length-additive expression for some $d\in W$ by [BB05, Proposition 3.1.3].
+That is, $a={{}_{I}w}=w_{0}(I)$ in the left parabolic decomposition of $w$.
+For the other direction, if $a=w_{0}(I)$, then $I\subseteq\mathcal{D}_{L}(w)$.
+Hence, (4) and (5) are equivalent.
+Finally, the equivalence of (2) and (5) is immediate. ∎
+Lemma 4.4.
+$\operatorname{Heart}_{I}(X_{w})$ is $U_{\bar{a}}$-stable for the usual action.
+Proof.
+We have that
+$U_{\bar{a}}\operatorname{Heart}_{I}(X_{w})$ $=U_{\bar{a}}aU_{d}a^{-1}wB=aa^{-1}U_{\bar{a}}aU_{d}dB=a\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{a^{-1}(\alpha)}U_{d}dB$
+$=aU_{d}dB=aU_{d}a^{-1}adB=\operatorname{Heart}_{I}(X_{w})$
+where the third equality follows by [J03, Part II, 1.4(5)], and the fourth equality is (5) combined with the fact that $\alpha\notin{\mathcal{I}}(a)$ implies $U_{a^{-1}(\alpha)}\leq B$. ∎
 Lemma 4.5.
@@
   1. (1)
-    $uh_{1}\notin\operatorname{Heart}_{I}(X_{w})$ for all $u\in U_{a}$ with $u\neq e$.
+    $u_{2}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$ for all $u_{2}\in U_{a}$ with $u_{2}\neq e$.
   2. (2)
-    $tx\notin\operatorname{Heart}_{I}(X_{w})$ for all $t\in T$.
+    $tu_{1}x\notin\operatorname{Heart}_{I}(X_{w})$ for all $t\in T$ and $u_{1}\in U_{\bar{a}}$.
   3. (3)
-    Let $b=tu_{1}u_{2}\in B_{L_{I}}$ with $t\in T$, $u_{1}\in U_{a}$, and $u_{2}\in U_{\bar{a}}$. If $bh_{1}=h_{2}$, then $u_{1}=e$ and $th_{1}=h_{2}$.
+    Let $b=tu_{1}u_{2}\in B_{L_{I}}$ with $t\in T$, $u_{1}\in U_{\bar{a}}$, and $u_{2}\in U_{a}$. If $bh_{1}=h_{2}$, then $u_{2}=e$ and $tu_{1}h_{1}=h_{2}$.
 Proof.
-(1) Since $h_{1}\in\operatorname{Heart}_{I}(X_{w})$, $h_{1}=vwB$ for some $v\in V_{d}$. By Lemma 2.3, $uv\notin V_{d}$. Thus $uvwB\in X_{w}^{\circ}\setminus\operatorname{Heart}_{I}(X_{w})$ by (5).
-(2) The $T$-stability of $\operatorname{Heart}_{I}(X_{w})$, proved in Proposition  4.2, implies $tx\notin\operatorname{Heart}_{I}(X_{w})$.
-(3) By Lemma 4.3, $u_{2}h_{1}=h_{1}$.
-Suppose, for contradiction that $u_{1}\neq e$.
-Then (1) implies that $u_{1}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$.
-And thus, (2) implies $tu_{1}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$.
+(1) Since $h_{1}\in\operatorname{Heart}_{I}(X_{w})$, $h_{1}=vwB$ for some $v\in V_{d}$.
+By Lemma 2.3, $u_{2}v\notin V_{d}$.
+Thus $u_{2}vwB\in X_{w}^{\circ}\setminus\operatorname{Heart}_{I}(X_{w})$ by (5).
+(2) The $T$ and $U_{\bar{a}}$ stability of $\operatorname{Heart}_{I}(X_{w})$, proved in Proposition  4.2 and Lemma 4.4 respectively, imply $tu_{1}x\notin\operatorname{Heart}_{I}(X_{w})$.
+(3) Suppose, for contradiction that $u_{2}\neq e$.
+Then (1) implies that $u_{2}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$.
+And thus, (2) implies $tu_{1}u_{2}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$.
 This contradicts our assumption that $tu_{1}u_{2}h_{1}=h_{2}$.
-Thus $u_{1}=e$ and $th_{1}=bh_{1}=h_{2}$.
+Thus $u_{2}=e$ and $tu_{1}h_{1}=bh_{1}=h_{2}$.
 ∎
-Finally, we are able to describe the precise form of $(B_{L_{I}})_{x}$.
 Lemma 4.6.
-Let $x\in X_{w}^{\circ}$.
-Then $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$.
-And $(B_{L_{I}})_{x}=T_{h}\ltimes U_{\bar{a}}$ for any $h\in(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})$.
+Let $x\in X_{w}^{\circ}$. Then $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$.
 Proof.
@@
 Thus $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$.
-Let $h\in(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})$.
-It is immediate by Corollary 4.4 that $T_{h}\ltimes U_{\bar{a}}\leq(B_{L_{I}})_{h}$.
-On the other hand, let $b\in(B_{L_{I}})_{h}$ so that $bh=h$.
-We have $b=tu_{1}u_{2}$ with $t\in T$, $u_{1}\in U_{a}$, and $u_{2}\in U_{\bar{a}}$.
-Lemma 4.5(3) implies that $u_{1}=e$ and $th=h$.
-Hence $b\in T_{h}\ltimes U_{\bar{a}}$.
-We conclude that $(B_{L_{I}})_{h}=T_{h}\ltimes U_{\bar{a}}$.
-Our proof is complete since $(B_{L_{I}})_{x}=(B_{L_{I}})_{h}$.
 ∎
-We now define our bijection between $B_{L_{I}}$-orbits in $X_{w}^{\circ}$ and $T$-orbits in $\operatorname{Heart}_{I}(X_{w})$.
+We now define our surjection from $T$-orbits in $\operatorname{Heart}_{I}(X_{w})$ to $B_{L_{I}}$-orbits in $X_{w}^{\circ}$, which is a codimension preserving bijection in the case where $L_{I}$ acts on the Schubert variety .
 Theorem 4.7.
-The map $\beta:\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))\rightarrow\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\mapsto B_{L_{I}}x$ where $x$ is any point in $\Theta$ is a codimension-preserving bijection.
+The map $\beta:\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))\rightarrow\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\mapsto B_{L_{I}}x$ where $x$ is any point in $\Theta$ is a surjection.
+If $L_{I}$ acts on $X_{w}$, then $\beta$ is a codimension preserving bijection.
 That is,
@@
 We begin by noting that $T\leq B_{L_{I}}$ implies that this map is well-defined.
+Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$.
+By Lemma 4.6, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$.
+Thus $h\in\Theta$ for some $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ and so $\beta(\Theta)=\Xi$.
+Thus $\beta$ is surjective.
+For the remainder of this proof we assume that $L_{I}$ acts on $X_{w}$; that is, by Lemma 4.3, $U_{\bar{a}}$ is the trivial subgroup of $G$.
 Let $\Theta_{1},\Theta_{2}\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ with $x\in\Theta_{1}$ and $y\in\Theta_{2}$.
@@
 Thus $b_{2}^{-1}b_{1}x=y$.
-Lemma 4.5(3) implies there exists a $t\in T$ such that $tx=y$.
+Lemma 4.5(3) implies there exists a $t\in T$ and $u\in U_{\bar{a}}$ such that $tux=y$.
+Since $U_{\bar{a}}$ is the trivial subgroup, this implies $tx=y$.
 Thus $\Theta_{1}=\Theta_{2}$ and $\beta$ is injective.
+Thus $\beta$ is a bijection.
 Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$.
-By Lemma 4.6, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$.
-Thus $h\in\Theta$ for some $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ and so $\beta(\Theta)=\Xi$.
-Thus $\beta$ is surjective.
-Continuing with this notation, we now consider the codimension of the orbits $\Xi$ and $\Theta$ in $X_{w}^{\circ}$ and $\operatorname{Heart}_{I}(X_{w})$, respectively.
-One has, by [B09, Proposition 1.11], that for any algebraic group $H$ and $H$-variety $X$, the orbit $H\cdot x$, $x\in X$, is a subvariety of $X$ of dimension $\dim H-\dim H_{x}$.
-Thus
-$\dim(\Xi)$ $=\dim(B_{L_{I}})-\dim(B_{L_{I}})_{x}$
-$=\dim(B_{L_{I}})-\dim(T_{h}\ltimes U_{\bar{a}})$
-$=(\ell(w_{0}(I))+\dim(T))-((\dim(T)-\dim(\Theta))+(\ell(w_{0}(I))-\ell(a)))$
+As above, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$ by Lemma 4.6.
+Thus $h\in\Theta$ for a unique $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ since $\beta$ is a bijection.
+We claim that
+$(B_{L_{I}})_{h}=T_{h}.$
+Clearly $T_{h}\subseteq(B_{L_{I}})_{h}$.
+For the other direction, suppose $bh=tu_{1}u_{2}h=h$ with $t\in T$, $u_{1}\in U_{\bar{a}}$, and $u_{2}\in U_{a}$.
+Then Lemma 4.5(3) implies $u_{2}=e$ and $tu_{1}h=h$.
+Since $U_{\bar{a}}$ is the trivial subgroup, $u_{1}=e$ and so $th=h$.
+We conclude $b=t\in T_{h}$.
+One has, by [B09, Proposition 1.11], that for any algebraic group $H$ and $H$-variety $X$, the orbit $H\cdot x$, $x\in X$, is a subvariety of $X$ of dimension $\dim H-\dim H_{x}$. Thus
+$\dim(\Xi)$ $=\dim(B_{L_{I}})-\dim((B_{L_{I}})_{h})$
+$=\dim(B_{L_{I}})-\dim(T_{h})$
+$=(\ell(w_{0}(I))+\dim(T))-((\dim(T)-\dim(\Theta))$
 $=\ell(a)+\dim(\Theta),$
-where the second equality is Lemma 4.6, and the third equality follows from the definition of $B_{L_{I}}$ and $U_{\bar{a}}$ as well as [B09, Proposition 1.11] applied to the $T$-orbit $\Theta$. Hence
+where the fourth equality is Lemma (4.3)(4). Hence
 $\dim(X_{w}^{\circ})-\dim(\Xi)$ $=(\ell(a)+\ell(d))-(\ell(a)+\dim(\Theta))$
@@
 Hence, Proposition 4.2 implies that the map $\mathfrak{T}:\mathcal{O}_{T}(X^{\circ}_{d})\rightarrow\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ given by $\Theta\mapsto Tax$, where $x$ is any point in $\Theta$, is a codimension preserving bijection.
-The map $\mathfrak{O}$ is precisely the composition of two codimension-preserving bijections, namely $\mathfrak{O}=\beta\circ\mathfrak{T}$.
-Hence $\mathfrak{O}$ is itself a codimension-preserving bijection.
+The map $\mathfrak{O}$ is precisely the composition of a surjection and a bijection, namely $\mathfrak{O}=\beta\circ\mathfrak{T}$.
+Hence $\mathfrak{O}$ is itself a surjection.
+In the case where $L_{I}$ acts on $X_{w}$, $\mathfrak{O}$ is the composition of two codimension-preserving bijections, and hence is a codimension preserving bijection.
 ∎
-Remark.
-For the remainder of this paper, we abuse notation slightly by writing $C_{B_{L_{I}}}(X^{\circ}_{w})$ and $C_{B_{L_{I}}}(X_{w})$ for the minimum codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ and $X_{w}$, respectively.
-This allows for discussing these orbit codimensions even when $L_{I}$ does not act on $X^{\circ}_{w}$ or $X_{w}$.
 Corollary 4.8.
-The minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ equals the torus complexity of $X^{\circ}_{d}$ for the usual torus action.
-That is, $c_{B_{L_{I}}}(X^{\circ}_{w})=c_{T}(X^{\circ}_{d})$.
+If $L_{I}$ acts on $X_{w}$, the minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ equals the torus complexity of $X^{\circ}_{d}$ for the usual torus action.
+That is, $c_{L_{I}}(X^{\circ}_{w})=c_{T}(X^{\circ}_{d})$.
 4.2. Orbits in the Schubert Variety
@@
 For any $u\in W$, let $u={{}_{I}u}\prescript{I}{}{u}$ be the left parabolic decomposition of $u$ with respect to $I$.
-The goal of this section is to use Theorem 4.7 to prove Corollary 1.6 and Theorem 1.5.
-Proof of Corollary 1.6. The Bruhat decomposition tells us that the Schubert variety $X_{w}$ is the disjoint union of the Schubert cells for $u\leq w$,
-$X_{w}=\displaystyle\bigsqcup_{u\leq w}X_{u}^{\circ}.$
-and hence
-(6) $\mathcal{O}_{B_{L_{I}}}(X_{w})=\bigsqcup_{u\leq w}\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u}).$
-Theorem 4.7 implies there is a bijection between $\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})$ and $\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})\times\{u\}$ for all $u\leq w$, which completes our proof.
-∎
-Remark.
-In Corollary 1.6 we decorate each of the sets $\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})$ by taking it’s product with $\{u\}$.
-This is due to the fact that for $u,v\leq w$ with $u\neq v$ it is possible that $\prescript{I}{}{u}=\prescript{I}{}{v}$.
-And we desire the union in the statement of Corollary 1.6 to be disjoint.
-Our next order of business is to apply Corollary 4.8 towards a formula for the minimal codimension of a $B_{L_{I}}$-orbit in any Schubert variety $X_{w}$.
+Our goal is to extend Corollary 4.8 to a formula for the Levi complexity of a Schubert variety.
+To do this, we will need to lower bound the codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{u}$ for $u\leq w$.
+The added complication in this case is that $L_{I}$ may not act on $X_{u}$, though of course $B_{L_{I}}$ acts, and so $U_{\bar{a}}$ will not be the trivial subgroup of $G$.
+For a subvariety $X$ of a variety $Y$, we write $\operatorname{codim}_{Y}(X)$ for the codimension of $X$ in $Y$.
 Proposition 4.9.
-For $w\in W$, the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$ is given by
-$c_{B_{L_{I}}}(X_{w})=\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right).$
+Let $u\leq w\in W$. Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})$. Let $h\in\Xi\cap\operatorname{Heart}_{I}(X_{u})$ and $\Theta=T\cdot h$. Then
+$\operatorname{codim}_{X^{\circ}_{u}}(\Xi)$ $\geq\ell(u)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I))$
+and
+$\operatorname{codim}_{X_{w}}(\Xi)\geq\ell(w)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I)).$
 Proof.
-For a $u\leq w$, the minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{u}$ is $\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})$ by Corollary 4.8 and Corollary 1.3.
-The codimension of $X^{\circ}_{u}$ in $X_{w}$ is $\ell(w)-\ell(u)$.
-These two preceding sentences combine with (6) to imply
-$c_{B_{L_{I}}}(X_{w})=$ $\min_{u\leq w}\left(\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})+\ell(w)-\ell(u)\right)$
-$=$ $\ell(w)+\min_{u\leq w}\left(-\ell(_{I}u)-\operatorname{supp}(\prescript{I}{}{u})\right)$
-$=$ $\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right),$
-where the second inequality is the length-additivity of the left parabolic decomposition. ∎
-The following example illustrates that this value is not, in general, well-behaved.
-Example 4.10.
-Let $w=56781234\in S_{8}$ written in one-line notation.
-Then $w$ is already in $\prescript{I}{}{W}$ since $\mathcal{D}_{L}(w)\cap I$ is the empty set.
-Therefore, $\ell(_{I}w)+\operatorname{supp}(\prescript{I}{}{w})=0+7=7$.
-Now if we apply $s_{4}$ to $w$ and get $u=56718234$, then $\prescript{I}{}{u}=25681347$, and ${}_{I}u$ has length $4$.
-Therefore, $\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})=4+7=11$.
-This is not the optimal choice for $u$: if we take $u=53218764\leq w$, then $\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})=6+7=13$.
-However, the formula in Proposition 4.9 can be simplified in restricted cases.
-Lemma 4.11.
-Let $w\in W$ such that $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$, i.e.
-$w$ admits a left Billey-Postnikov decomposition with respect to $I$ (see [RS17, Proposition 4.2]).
-Then $w$ simultaneously maximizes both $\ell(_{I}u)$ and $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$.
-Thus
-$c_{B_{L_{I}}}(X_{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
-Proof.
-For any $w$, if $u\leq w$, then $\prescript{I}{}{u}\leq\prescript{I}{}{w}$.
-Thus $w$ always maximizes $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$.
-Hence, if $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$, then $w$ simultaneously maximizes both $\ell(_{I}u)$ and $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$.
-Applying this to Proposition 4.9 yields
-$c_{B_{L_{I}}}(X_{w})=$ $\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right)$
-$=$ $\ell(w)-\ell(_{I}w)-\operatorname{supp}(\prescript{I}{}{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
-The hypothesis of Lemma 4.11 is satisfied when $L_{I}$ acts $X_{w}$.
-This yields one of our main results, Theorem 1.5: a simple, type independent formula for the Levi complexity of a Schubert variety.
+The second claimed inequality follows from the first combined with the fact that the codimension of $X^{\circ}_{u}$ in $X_{w}$ is $\ell(w)-\ell(u)$.
+Hence it remains to prove the first inequality.
+By [B09, Proposition 1.11] $\dim(\Xi)=\dim(B_{L_{I}})-\dim(B_{h})\leq\dim(B_{L_{I}})-\dim(T_{h})$.
+Multiplying this inequality by $-1$ and adding $\ell(u)$ yields
+(6) $\begin{split}\ell(u)-\dim(\Xi)&\geq\ell(u)-\dim(B_{L_{I}})+\dim(T_{h})\\ &=\ell({{}_{I}u})+\ell(\prescript{I}{}{u})-(\dim(T)+\ell(w_{0}(I)))+\dim(T_{h})\\ &=\ell(\prescript{I}{}{u})-(\dim(T)-\dim(T_{h}))-(\ell(w_{0}(I))-\ell({{}_{I}u}))\\ &=\ell(\prescript{I}{}{u})-\dim(\Theta)-(\ell(w_{0}(I))-\ell({{}_{I}u})),\end{split}$
+where the first equality is the length additivity of $u={{}_{I}u}\prescript{I}{}{u}$ and the second equality is [B09, Proposition 1.11].
+Proposition 4.2 implies that $\operatorname{Heart}_{I}(X_{u})$ is $T$-equivariantly isomorphic to $X^{\circ}_{\prescript{I}{}{u}}$.
+Hence, $\dim(\operatorname{Heart}_{I}(X_{u}))=\ell(\prescript{I}{}{u})$ and
+(7) $\ell(\prescript{I}{}{u})-\dim(\Theta)=\operatorname{codim}_{\operatorname{Heart}_{I}(X_{u})}(\Theta)\geq c_{T}(X^{\circ}_{\prescript{I}{}{u}})\geq c_{T}(X_{\prescript{I}{}{u}}).$
+Thus,
+$\operatorname{codim}_{X^{\circ}_{u}}(\Xi)$ $=\dim(X^{\circ}_{u})-\dim(\Xi)=\ell(u)-\dim(\Xi)$
+$\geq\ell(\prescript{I}{}{u})-\dim(\Theta)-(\ell(w_{0}(I))-\ell({{}_{I}u}))\geq c_{T}(X_{\prescript{I}{}{u}})-(\ell(w_{0}(I))-\ell({{}_{I}u}))$
+$=\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})-(\ell(w_{0}(I))-\ell({{}_{I}u}))=\ell(u)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I)),$
+where the first inequality is (6), the second inequality is (7), the third equality is Corollary 1.3, and the final inequality is the length additivity of $u={{}_{I}u}\prescript{I}{}{u}$. ∎
 We are now ready to prove Theorem 1.5.
 Proof of Theorem 1.5.
-The equivalence of $L_{I}$ acting on $X_{w}$ and $I\subseteq\mathcal{D}_{L}(w)$ is [BL00, Lemma 8.2.3].
-If $I\subseteq{\mathcal{D}}_{L}(w)$, then ${\mathcal{I}}(w)$ contains all the positive roots in the root subsystem generated by $I$.
-Thus $w=w_{0}(I)d$ is a length-additive expression for some $d\in W$ by [BB05, Proposition 3.1.3].
-That is, ${}_{I}w=w_{0}(I)$ in the left parabolic decomposition of $w$.
-Thus, it must be the case then that $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$.
-Hence Lemma 4.11 implies the theorem statement.
+The equivalence of $L_{I}$ acting on $X_{w}$ and $I\subseteq\mathcal{D}_{L}(w)$ is Lemma 4.3.
+The Bruhat decomposition tells us that the Schubert variety $X_{w}$ is the disjoint union of the Schubert cells for $u\leq w$, $X_{w}=\bigsqcup_{u\leq w}X_{u}^{\circ}$, and hence
+$\mathcal{O}_{B_{L_{I}}}(X_{w})=\bigsqcup_{u\leq w}\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u}).$
+Thus
+$c_{B_{L_{I}}}(X_{w})$ $=\min_{\begin{subarray}{c}u\leq w\\ \Xi\in\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})\end{subarray}}\operatorname{codim}_{X_{w}}(\Xi)$
+$\geq\min_{u\leq w}\left(\ell(w)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I))\right)$
+$=\ell(w)-\ell(w_{0}(I))+\min_{u\leq w}\left(-\operatorname{supp}(\prescript{I}{}{u})\right)$
+$=\ell(\prescript{I}{}{w})-\max_{u\leq w}\left(\operatorname{supp}(\prescript{I}{}{u})\right)$
+$=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}),$
+where the inequality is Proposition 4.9, the third equality follows from Lemma 4.3, and the final equality follows from the fact that $u\leq w$ implies $\prescript{I}{}{u}\leq\prescript{I}{}{w}$.
+Corollary 4.8 and Corollary 1.3 imply that this lower bound on $c_{B_{L_{I}}}(X_{w})$ is in fact achieved.
+We conclude that $c_{B_{L_{I}}}(X_{w})=\ell(\prescript{I}{}{w})+\operatorname{supp}(\prescript{I}{}{w})$.
 ∎